### v2

# Near-Optimal Regret for Efficient Stochastic Combinatorial Semi-Bandits

Abstract
The combinatorial multi-armed bandit (CMAB) is a cornerstone of sequential decision-making framework, dominated by two algorithmic families: UCB-based and adversarial methods such as follow the regularized leader (FTRL) and online mirror descent (OMD). However, prominent UCB-based approaches like CUCB suffer from additional regret factor $\log T$ that is detrimental over long horizons, while adversarial methods such as EXP3.M and HYBRID impose significant computational overhead. To resolve this trade-off, we introduce the Combinatorial Minimax Optimal Strategy in the Stochastic setting (CMOSS). CMOSS is a computationally efficient algorithm that achieves an instance-independent regret of $O\big((\log k)\sqrt{kmT}\big)$ when $k\leq\frac{m}{2}$ and $O\big((m-k)\sqrt{\log k\log(m-k)T}\big)$ when $k>\frac{m}{2}$ under semi-bandit feedback, where $m$ is the number of arms and $k$ is the maximum cardinality of a feasible action. Crucially, this result eliminates the dependency on $\log T$ and matches the established lower bounds of $\Omega\big(\sqrt{kmT}\big)$ when $k\leq\frac{m}{2}$ and $\Omega\big((m-k)\sqrt{\log(\frac{m}{m-k})T}\big)$ when $k>\frac{m}{2}$ up to logarithmic terms of $k$ and $m$. We then extend our analysis to show that CMOSS is also applicable to cascading feedback. Experiments on synthetic and real-world datasets validate that CMOSS consistently outperforms benchmark algorithms in both regret and runtime efficiency.
1 Introduction
The stochastic multi-armed bandit (MAB) is a fundamental $T$-round sequential decision-making game that has been widely studied in the past two decades (Audibert and Bubeck 2009; Auer et al. 1995, 2002; Robbins 1952). In this game, one player faces $m$ arms. In each round $t\in[T]$, the player chooses an arm $A_{t}$ among the $m$ arms and gets a reward drawn from a predetermined but unknown reward distribution with mean $\mu_{A_{t}}\in[0,1]$. The arm with the largest reward mean $\mu_{*}$ is called the best arm. The total expected regret is defined as $\mathbf{E}\left[R(T)\right]=\mathbf{E}\left[\sum_{t=1}^{T}\left(\mu_{*}-\mu_{A_{t}}\right)\right],$ where the expectation is over the randomness from the player’s strategy and the random reward. And the objective is to minimize the minimax regret, namely the regret of the best algorithm against the worst input.
The Upper Confidence Bounds (UCB) algorithm (Auer, Cesa-Bianchi, and Fischer 2002) is a well-known approach for addressing the stochastic MAB problem, achieving an expected regret of $O\left(\sqrt{mT\log T}\right)$. Building on this foundation, its variant, the Minimax Optimal Strategy in the Stochastic case (MOSS) algorithm (Audibert and Bubeck 2009; Degenne and Perchet 2016; Ménard and Garivier 2017), improves the regret bound to $O\left(\sqrt{mT}\right)$ by refining the upper confidence bounds and employing a distinct regret analysis. These results align with the minimax lower bound of $\Omega\left(\sqrt{mT}\right)$ established in (Auer et al. 2002), demonstrating their optimality in the stochastic setting. In parallel, the EXP3 algorithm (Auer et al. 1995; Littlestone and Warmuth 1994) provides a solution for adversarial environments. EXP3 is a classic implementation of the follow the regularized leader (FTRL) framework using a negative entropy regularizer (Shalev-Shwartz et al. 2012). This approach, also known through the dual perspective of online mirror descent (OMD), achieves a regret of $O\left(\sqrt{mT\log m}\right)$ in both adversarial and stochastic regimes, notably avoiding dependence on $\log T$.
In many real-world applications, however, decision-making problems often extend beyond the simple MAB framework to involve combinatorial actions among multiple arms. For instance, in online advertising (Kveton et al. 2015a), an advertiser could select up to $k$ web pages to display ads, constrained by a budget. Each user visits a subset of pages, and the advertiser must learn the unknown click-through probabilities to maximize clicks. Similarly, in viral marketing (Gai, Krishnamachari, and Jain 2012; Tang et al. 2017), a marketer selects seed nodes in a social network to initiate information cascades, learning influence probabilities to maximize the reach of campaigns. These scenarios highlight the need for algorithms that efficiently address the combinatorial MAB (CMAB) problem.
While a number of works (Combes et al. 2015; Kveton et al. 2015c; Chen et al. 2016; Wang and Chen 2017; Liu et al. 2022, 2024, 2025) have focused on the CMAB problem, existing algorithms either suffer from additional $\log T$ terms in their regret bounds or incur high computational costs in practice. Among them, the Combinatorial UCB (CUCB) algorithm (Chen, Wang, and Yuan 2013; Kveton et al. 2015c; Wang and Chen 2017; Liu et al. 2022) is one of the most prominent and widely adopted approaches. The CUCB algorithm extends the classic UCB framework by selecting the action (super arm) from a feasible set that solves a combinatorial optimization problem defined over the upper confidence bounds of the base arms. Kveton et al. (Kveton et al. 2015c) establish an upper regret bound of $O\left(\sqrt{kmT\log T}\right)$ for CUCB, which matches the lower bound of $\Omega\left(\sqrt{kmT}\right)$ only up to a $\sqrt{\log T}$ factor when $k\leq\frac{m}{2}$.
A separate line of work (Cesa-Bianchi and Lugosi 2012; Combes et al. 2015; Uchiya, Nakamura, and Kudo 2010; Zimmert, Luo, and Wei 2019) focuses on adapting algorithms derived from FTRL and OMD frameworks that could be applied to CMAB setting. The best-known result in this line (Zimmert, Luo, and Wei 2019) achieves a regret of $O\left(\sqrt{kmT}\right)$ when $k\leq\frac{m}{2}$ and $O\left((m-k)\sqrt{\log\left(\frac{m}{m-k}\right)T}\right)$ when $k>\frac{m}{2}$, successfully removing the $\log T$ term. However, these adversarially designed algorithms are often computationally expensive, limiting their practical utility. This gap—between computational efficiency and regret optimality—has remained unresolved for over a decade, highlighting the critical need for “MOSS-like” algorithms in CMAB that achieve minimax-optimal regret without $\log T$ dependence.
1.1 Our results
To address the above limitation, we study the stochastic combinatorial semi-bandit problem and propose the CMOSS algorithm. Our contribution is to show that CMOSS is not only computationally efficient but also has a regret of $O\big((\log k)\sqrt{kmT}\big)$ when $k\leq\frac{m}{2}$ and $O\big((m-k)\sqrt{\log k\log(m-k)T}\big)$ when $k>\frac{m}{2}$. We also extend the setting to cascading feedback and compare our result with CUCB, since the other three algorithms can not be applied to the scenarios where arms in $A_{t}$ are probabilistically triggered. The results are summarized in Table 1.
Table 1: Regret bounds and empirical computational cost per-round for algorithms on the CMAB problem.
$\star$ COMBEXP is from (Combes et al. 2015). CUCB is from (Chen, Wang, and Yuan 2013; Kveton et al. 2015c; Liu et al. 2024). EXP3.M is from (Uchiya, Nakamura, and Kudo 2010). HYBRID is from (Zimmert, Luo, and Wei 2019). The detailed regret bound of CMOSS is given in Theorem 1 and Theorem 5.
$\circ$ “None” means these methods can not deal with this case. $p^{*}$ denotes the minimal probability that an action could have all base arms observed over all time.
$\dagger$ We select $k=20,m=30$ for HYBRID and CMOSS when $k>\frac{m}{2}$ and $k=10,m=30$ for other rows with $T=100000$ separately for runtime measurement on a synthetic dataset (averaged over 10 runs) in the semi-bandit feedback setting. See Section 4.1 for dataset details.
$\ddagger$ HYBRID performs relatively bad in experiments simulating realistic conditions where arm means are in the range of $[0,0.1]$, although the regret bound is optimal. See Section 4.1 for dataset details.
[…]
COMBEXP $O\left(\sqrt{k^{3}mT\log m}\right)$ None Too large
CUCB $O\left(\sqrt{kmT\log T}\right)$ $O\left(\sqrt{kmT\log T}\right)$ 1.516
EXP3.M $O\left(\sqrt{kmT\log\left(\frac{m}{k}\right)}\right)$ None 10.429
[…]
$O\left((m-k)\sqrt{\log\left(\frac{m}{m-k}\right)T}\right)$, $k>\frac{m}{2}$ None 11.195
CMOSS $O\left((\log k)\sqrt{kmT}\right)$, $k\leq\frac{m}{2}$ $O\left(\frac{\log k}{p^{*}}\sqrt{kmT}\right)$ 1.737
$O\left((m-k)\sqrt{\log k\log(m-k)T}\right)$, $k>\frac{m}{2}$ $O\left(\frac{m-k}{p^{*}}\sqrt{\log k\log(m-k)T}\right)$ 3.067
The $\log T$ factor of many UCB-style algorithms is due to ensuring that the empirical mean of every arm must lie within its confidence interval. To achieve optimal instance-independent regret bound, the MOSS algorithm bypasses this requirement. MOSS operates by classifying arms and then directly bounding the expected plays of suboptimal arms with relatively large gaps, using modified confidence intervals to facilitate this analysis. However, this is not directly tractable in the CMAB setting. First, how to classify the suboptimal actions based on their gaps; second, how to bound the expected number of times each suboptimal action is pulled; and finally, how to tractably aggregate the regret contributions from a potentially exponential number of such actions.
Our technical contribution lies in bridging the gap between the MOSS-style analysis and the combinatorial structure of CMAB, which prior work such as CUCB does not address. To this end, we introduce new analytical tools that enable regret-optimal and computationally efficient learning. Our key innovations include:
  1. 1.
    Targeted action filtering. We identify suboptimal actions with gaps exceeding the worst-case underestimation error of the optimal action. Actions with gaps smaller than a chosen threshold $\Delta$ contribute at most $T\Delta$ to the regret, allowing us to focus on a tractable subset of actions.
  2. 2.
    Constraint decoupling via classification. In combinatorial settings, each action involves multiple arms, which complicates concentration analysis. To address this, we introduce a partitioning scheme that classifies suboptimal actions based on structural properties. This enables us to decouple constraints at the arm level and apply refined concentration bounds.
  3. 3.
    Regret decomposition. After the preparation of the first two points, we need to choose a proper way to calculate the regret to avoid redundant computations. Since each suboptimal arm may participate in an exponential number of actions, directly summing over all possibilities is infeasible. We instead decompose the total regret into per-arm contributions and use an integral-based approximation to tightly and efficiently estimate the overall impact. See Section 3.2 for more detailed analysis.
Furthermore, we show that CMOSS could also be extended to cascading feedback. More details are given in Section 3.3. We also run comparative experiments evaluating the performance of CUCB and CMOSS in Section C.4.
We also conduct comparative experiments involving CMOSS, CUCB, EXP3.M, and HYBRID on both synthetic and real-world datasets under semi-bandit feedback, including ablation studies. Detailed descriptions of CUCB, EXP3.M, and HYBRID are provided in Appendix B for completeness. The results demonstrate that, under noisy conditions and after sufficient exploration rounds, CMOSS incurs slightly higher computational cost than CUCB while achieving notably lower cumulative regret. Compared to EXP3.M and HYBRID, CMOSS attains better cumulative regret with significantly reduced computational cost. This performance profile highlights CMOSS’s superior balance between exploration and exploitation, underscoring its effectiveness in such environments. Further details are provided in Section 4.
1.2 Related work
Multi-Armed Bandits (MAB). The MAB problem was first introduced in (Robbins 1952). Audibert et al. (Audibert and Bubeck 2009) later established an optimal regret bound of $\Theta(\sqrt{mT})$ for both stochastic and adversarial settings. Our work focuses on the stochastic setting, where the UCB algorithm (Auer, Cesa-Bianchi, and Fischer 2002) plays a pivotal role and serves as the foundation for many CMAB studies. To eliminate the $\log T$ factor, the MOSS algorithm was proposed in (Audibert and Bubeck 2009), with an anytime modification introduced in (Degenne and Perchet 2016), achieving a regret of $O\left(\sqrt{mT}\right)$. In contrast, works such as (Auer et al. 1995; Littlestone and Warmuth 1994) targeted the adversarial setting and proposed the EXP3 algorithm, which also eliminates the $\log T$ factor and achieves a regret of $O\left(\sqrt{mT\log m}\right)$. However, EXP3 incurs significantly higher computational overhead compared to UCB-based methods such as MOSS. Our work builds primarily on MOSS and extends it to the combinatorial setting.
Stochastic Combinatorial Multi-Armed Bandits (CMAB). The stochastic CMAB problem has been extensively studied in the literature (Chen et al. 2016; Chen, Wang, and Yuan 2013; Combes et al. 2015; Kveton et al. 2015a, c; Wang and Chen 2017; Liu et al. 2022, 2023). The seminal work by Gai et al. (Gai, Krishnamachari, and Jain 2012) pioneered the study of stochastic CMAB with semi-bandit feedback, which inspired algorithmic advancements and improved regret bound across various settings (Kveton et al. 2015a, 2014; Tang et al. 2017). For the instance-dependent regret, Wang et al. (Wang and Chen 2018) proposed the combinatorial Thompson sampling (CTS) algorithm with an offline oracle, achieving a regret of $O\left(\frac{m\log T}{\Delta_{\min}}\right)$, where $\Delta_{min}$ is the minimum gap between the expected reward of the optimal action and any suboptimal action. The bound essentially matches asymptotically on $T$. Thus in this paper, we focus on the instance-independent regret.
For the instance-independent regret, one line of works focused on the adversarial setting and designed algorithms based on the principle of OMD and FTRL. Several studies (Cesa-Bianchi and Lugosi 2012; Dani, Kakade, and Hayes 2007; Uchiya, Nakamura, and Kudo 2010; Zimmert, Luo, and Wei 2019) proposed enhancements such as COMBEXP, EXP3.M, with the best (Zimmert, Luo, and Wei 2019) achieving a regret of $O\left(\sqrt{kmT}\right)$ when $k\leq\frac{m}{2}$ and $O\left((m-k)\sqrt{\log\left(\frac{m}{m-k}\right)T}\right)$ when $k>\frac{m}{2}$. The latter case matches the lower bound of $\Omega\left((m-k)\sqrt{\log\left(\frac{m}{m-k}\right)T}\right)$. However, these methods suffer from high computational overhead. Another line of works (Chen, Wang, and Yuan 2013; Liu et al. 2022; Wang and Chen 2017, 2018; Liu et al. 2023) used UCB-based methods. Chen et al. (Chen, Wang, and Yuan 2013) proposed the CUCB algorithm and analyzed the regret of CUCB with an approximation oracle. Then Kveton et al. (Kveton et al. 2015c) derived an upper bound of $O\left(\sqrt{kmT\log T}\right)$ for CUCB and established a lower bound of $\Omega\left(\sqrt{kmT}\right)$ when $k\leq\frac{m}{2}$. To address the degradation caused by the $\log T$ term over long horizons, we refine the upper confidence bound and propose the CMOSS algorithm. It achieves strong performance while maintaining low computational costs, even after removing the $\log T$ term.
2 Preliminaries
We begin by introducing the general framework of the stochastic CMAB and establishing notations used throughout the paper. The CMAB problem is modeled as a sequential decision-making game between a learning agent and an environment. The environment consists of $m$ base arms $\mathtt{arm}_{1},\ldots,\mathtt{arm}_{m}$, each associated with a random reward $X_{i}\sim D_{i}$ supported on $[0,1]$ and having mean $\mu_{i}$. The distributions $\{D_{i}\}_{i=1}^{m}$ are fixed by the environment before the game begins and are drawn from a known class $\mathcal{D}$. The learner knows $\mathcal{D}$ but not the specific instantiation of each $D_{i}$. In each round, the rewards of these arms may be correlated with each other.
The action space $\mathcal{A}$ consists of all subsets of $[m]$ such that $\forall A\in\mathcal{A}$, $\left|A\right|\leq k$. The $T$-round decision game starts as follows: in each round $t\in[T]$, the player selects an action $A_{t}$ from $\mathcal{A}$ based on the feedback history, and the environment draws from every distribution $D_{i}$ a sample $X^{(t)}_{i}$. Let $\bm{\mu}$ represents the vector of the mean rewards of all arms. After the action $A_{t}$ is played, under semi-bandit feedback, the values of $X^{(t)}_{i}$ for all $i\in A_{t}$ are observed as the feedback to the player. Under cascading feedback, the arms in $A_{t}$ are probabilistically triggered and only the arms in the triggered set $\tau_{t}$ could be observed. In both cases the player receives $r(A_{t},\bm{\mu})$ as the reward in round $t$.
The player’s objective is to minimize the expected difference between the cumulative reward of the best action $A^{*}$ and the player’s own cumulative reward. That is, the player aims to design an algorithm to minimize the expected regret
$\mathbf{E}\left[R(T)\right]$ $=\mathbf{E}\left[\sum_{t=1}^{T}\left(r(A^{*},\bm{\mu})-r(A_{t},\bm{\mu})\right)\right].$
Notations. $\mathtt{U}(a,b)$ represents the uniform distribution on $(a,b)$. Denote $\Delta_{A_{t}}=r(A^{*},\bm{\mu})-r(A_{t},\bm{\mu})$ as the gap of $A_{t}$. $\mathbb{1}\left[S\right]$ is the indicator variable of whether event $S$ happens. For every $i\in[m]$, we use a random variable $T_{i,t}:=\sum_{s=1}^{t}\mathbb{1}\left[i\in A_{s}\right]$ to denote the number of times $\mathtt{arm}_{i}$ has been pulled when the $t$-th round ends. For two sets $A$ and $B$, $A\setminus B=\left\{a\;\middle|\;a\in A\land a\notin B\right\}$. And $\log$ represents the logarithm with base $2$. $\lfloor x\rfloor$ represents the largest integer that is no more than $x$. $\ln^{+}(x)=\ln\max\left\{1,x\right\}$.
3 CMOSS Algorithm and Its Regret Analysis
In this section, we first define CMOSS and present the theorem for its upper bound in Section 3.1. Next, in Section 3.2, we focus on the case of semi-bandit feedback, outline the proof idea and highlight key lemmas. Finally, in Section 3.3, we extend our analysis to cascading feedback.
3.1 The CMOSS algorithm
Algorithm 1 Combinatorial Minimax Optimal Strategy in the Stochastic setting (CMOSS)
Input:Number of base arms $m$, cardinality $k$, parameter $\delta$.
1: For each arm $i$, $T_{i}\leftarrow 0$ $\triangleright$ maintain the total number of times arm $i$ is played so far
2: For each arm $i$, $\hat{\mu}_{i}\leftarrow 1$ $\triangleright$ maintain the empirical mean of $\mathtt{arm}_{i}$
3: for $t=1,2,3,\ldots$ do
4:   For each arm $i\in[m]$, $\rho_{i}\leftarrow\sqrt{\frac{1}{T_{i}}\ln^{+}\left(\frac{1}{\delta T_{i}}\right)}$ $\triangleright$ the confidence radius, $\rho_{i}=+\infty$ if $T_{i}=0$
5:   For each arm $i\in[m]$, $\bar{\mu}_{i}\leftarrow\min\left\{\hat{\mu}_{i}+\rho_{i},1\right\},\bar{}\bm{\mu}\leftarrow(\bar{\mu}_{1},\cdots,\bar{\mu}_{m})$
6:   $A_{t}\leftarrow\mathop{\arg\max}_{A\in\mathcal{A}}r(A,\bar{}\bm{\mu})$
7:   Pull $A_{t}$ of base arms and obtain the triggered set $\tau_{t}$. Observe feedback $X_{i}^{(t)}$ for all $i\in\tau_{t}$
8:   For every $i\in\tau_{t}$, update $T_{i}$ and $\hat{\mu}_{i}$: $T_{i}\leftarrow T_{i}+1$, $\hat{\mu}_{i}\leftarrow\hat{\mu}_{i}+(X_{i}^{(t)}-\hat{\mu}_{i})/T_{i}$
The CMOSS algorithm maintains the empirical estimate $\hat{\mu}_{i}$ for the true mean $\mu_{i}$ for all the $m$ arms, then selects the action by solving the combinatorial problem of maximizing $r(A,\bar{}\bm{\mu})$, where each $\bar{\mu}_{i}$ optimistically estimates $\mu_{i}$ by a confidence interval $\rho_{i}$. The prominent difference between CMOSS and CUCB is that the radius $\rho_{i}=\sqrt{\frac{1}{T_{i}}\ln^{+}\left(\frac{1}{\delta T_{i}}\right)}$ but not $\sqrt{\frac{3\ln t}{2T_{i}}}$, where $\delta$ is a parameter to be optimized during the proof. In this case, the probability that $\hat{\mu}_{i}$ lies in the confidence region $(\mu_{i}-\rho_{i},\mu_{i}+\rho_{i})$ in every round $t$ could be much less than $\Theta\left(\frac{1}{t^{2}}\right)$. Thus we can not neglect the regret caused by $\hat{\mu}_{i}$ falling outside $(\mu_{i}-\rho_{i},\mu_{i}+\rho_{i})$. Therefore we need to bound the regret in the average sense. We have the following theorem:
Theorem 1.
For any stochastic combinatorial semi-bandit feedback instance with $m$ arms, time horizon $T$, and the maximum number of arms that could be chosen in a round is $k$, then when $k\leq\frac{m}{2}$, the expected regret of Algorithm 1 is $O\left((\log k)\sqrt{kmT}\right)$ and when $k>\frac{m}{2}$, the expected regret is $O(\left(m-k\right)\sqrt{\log k\log(m-k)T})$.
Theorem 1 eliminates the $\sqrt{\log T}$ factor in the upper bound of CUCB. Unlike the way EXP3.M and HYBRID do, CMOSS also retains the low computation cost, introducing only an additional logarithmic term of $k$ and $m$. And when k=mck=m-c ($c$ is a constant), our upper bound strictly matches the lower bound, which is $\Theta\left(c\sqrt{T\log m}\right)$.
3.2 Regret analysis for semi-bandit feedback
In this case $r(A_{t},\bm{\mu})=\sum_{i\in A_{t}}\mu_{i}$. Since $\mathcal{A}$ consists of all subsets of $[m]$ with cardinality less than $k$, and the reward function is monotonically increasing, then $A^{*}$ consists of $k$ arms corresponding to the largest values of $\mu_{i}$. $A_{t}$ is the set of $k$ arms corresponding to the largest values of $\bar{\mu}_{i}$ in round $t$. Since $\Delta_{A_{t}}=\sum_{j\in A^{*}}\mu_{j}-\sum_{i\in A_{t}}\mu_{i}$, we have
$\mathbf{E}\left[R(T)\right]$ $=\mathbf{E}\left[\sum_{t=1}^{T}\left(\sum_{j\in A^{*}}\mu_{j}-\sum_{i\in A_{t}}\mu_{i}\right)\right]$
$=\mathbf{E}\left[\sum_{t=1}^{T}\Delta_{A_{t}}\right].$
Denote $\hat{\mu}_{i,t},\rho_{i,t}$ as the value of $\hat{\mu}_{i}$ and $\rho_{i}$ when the $t$-th round ends. There are three pivotal challenges that are solved in our analysis:
  1. 1.
    Extracting the key actions to consider. We focus on bounding the regret for actions whose gap is larger than the summation of $\mu_{j}-\min\limits_{t\leq T}\left(\hat{\mu}_{j,t}+\rho_{j,t}\right)$ across all $\mathtt{arm}_{j}$ in $A^{*}$, since there are $k$ arms in $A^{*}$ and $A_{t}$ at each round $t$;
  2. 2.
    Decoupling constraints across suboptimal action $A_{t}$ into individual arms. We classify the action set into well-defined categories (Lemma 2) and quantitatively characterize the necessary condition that must be satisfied by the suboptimal arms belong to $A_{t}$ in each category Lemma 3);
  3. 3.
    Calculating the regret. We decompose the overall regret into contributions from each suboptimal arm. To manage the exponential number of actions for every suboptimal arm, we construct an integral to bound their summation, which effectively reduces redundant computations.
In this section we only present the case where $k\leq\frac{m}{2}$. All proofs of the lemmas in this section and the other case are deferred to Section A.1. Here $\delta$ is optimized to be $\frac{m(\log k)^{2}}{kT}$. Define $\zeta_{j}=\max\left(0,\mu_{j}-\min\limits_{t\leq T}\left(\hat{\mu}_{j,t}+\rho_{j,t}\right)\right)$ for every $j\in A^{*}$. We decompose the overall regret as follows:
$\phantom{{}={}}\sum_{t=1}^{T}\mathbf{E}\left[\Delta_{A_{t}}\right]$
$\leq\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[\Delta_{A_{t}}\leq(\log k)\sqrt{\frac{ekm}{T}}\right]\Delta_{A_{t}}\right]+$
$\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[\Delta_{A_{t}}\leq 8\sum_{j\in A^{*}}\zeta_{j}\right]\Delta_{A_{t}}\right]+$
$\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[\Delta_{A_{t}}>\max\left\{8\sum_{j\in A^{*}}\zeta_{j},(\log k)\sqrt{\frac{ekm}{T}}\right\}\right]\Delta_{A_{t}}\right]$
$\leq\sqrt{ekmT}+8\sum_{j\in A^{*}}\mathbf{E}\left[\zeta_{j}\right]\cdot T+$
$\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[\Delta_{A_{t}}>\max\left\{8\sum_{j\in A^{*}}\zeta_{j},(\log k)\sqrt{\frac{ekm}{T}}\right\}\right]\Delta_{A_{t}}\right].$ (1)
From Lemma 8 in Section A.1, which is proved by using concentration inequalities, the second term could be bounded as
$8\sum_{j\in A^{*}}\mathbf{E}\left[\zeta_{j}\right]\cdot T\leq 32k\sqrt{\delta}\cdot T=32(\log k)\sqrt{kmT}.$
Now we begin to bound the third term $\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{B}\right]\Delta_{A_{t}}\right]$, where
$\mathcal{A}_{B}=\left\{A\in\mathcal{A}\;\middle|\;\Delta_{A}>\max\left\{8\sum_{j\in A^{*}}\zeta_{j},(\log k)\sqrt{\frac{ekm}{T}}\right\}\right\}.$
This part consists of challenges 2 and 3. For challenge 2, we first show that when Algorithm 1 chooses $A_{t}$, for every arm ($\mathtt{arm}_{i}$ e.g.) in $A_{t}\setminus A^{*}$, we have
$\hat{\mu}_{i,t}+\rho_{i,t}$ $\geq\frac{1}{\left|A_{t}\setminus A^{*}\right|}\left(\sum_{j\in A^{*}\setminus A_{t}}\mu_{j}-\sum_{j\in A^{*}}\zeta_{j}\right)$
$\geq\frac{1}{\left|A_{t}\setminus A^{*}\right|}\left(\sum_{i\in A_{t}\setminus A^{*}}\mu_{i}+\frac{7\Delta_{A_{t}}}{8}\right),$
where the second inequality comes from $\Delta_{A_{t}}>8\sum_{j\in A^{*}}\zeta_{j}$. For every suboptimal action $A$, define $k_{A}=\left|A\setminus A^{*}\right|$. So $1\leq k_{A}\leq k$. We denote $\left(\mu^{A}_{i}\right)_{i\in[k_{A}]}$ as the mean rewards of the arms in $A\setminus A^{*}$ where $\mu^{A}_{k_{A}}\geq\mu^{A}_{k_{A}-1}\geq\cdots\geq\mu^{A}_{1}$ and $\overline{\mu^{A}}=\frac{\sum_{i\in[k_{A}]}\mu^{A}_{i}}{k_{A}}$. Define
$D_{A}=\left\{i\in[k_{A}]\;\middle|\;\mu^{A}_{i}\leq\overline{\mu^{A}}+\frac{3\Delta_{A}}{4k_{A}}\right\}$
as the set of arms in $A\setminus A^{*}$ whose mean reward is no more than the $\overline{\mu^{A}}+\frac{3\Delta_{A}}{4k_{A}}$. Then we have the following lemma:
Lemma 2.
For every suboptimal action $A\in\mathcal{A}$, if $\left|D_{A}\right|\leq\lfloor\frac{k_{A}}{2}\rfloor$, then there exists $1\leq\ell_{A}\leq\lfloor\log k_{A}\rfloor$ such that
$\overline{\mu^{A}}-\mu^{A}_{\lfloor\frac{k_{A}}{2^{\ell_{A}}}\rfloor}\geq\frac{\left(\sqrt{2}\right)^{\ell_{A}}}{16}\cdot\frac{\Delta_{A}}{k_{A}}.$
Lemma 2 reveals that every suboptimal action $A_{t}$ could be quantitatively characterized by a fundamental trade-off: it must contain either a few arms with mean rewards far from the action’s average mean, $\overline{\mu^{A}}$, or conversely, many arms with rewards clustered near it. Therefore before the algorithm starts, we could divide the action set $\mathcal{A}$ into $1+\lfloor\log k\rfloor$ exhaustive and mutually exclusive sets $\left(\mathcal{A}_{\ell}\right)_{0\leq\ell\leq\lfloor\log k\rfloor}$, where $\mathcal{A}_{0}=\left\{A\in\mathcal{A}\;\middle|\;\left|D_{A}\right|>\lfloor\frac{k_{A}}{2}\rfloor\right\}$, $\mathcal{A}_{\ell}=\left\{A\in\mathcal{A}\;\middle|\;\left|D_{A}\right|\leq\lfloor\frac{k_{A}}{2}\rfloor\land\ell_{A}=\ell\right\}$. If the choice of $\ell_{A}$ for $A$ is not unique, we choose $\ell_{A}$ to be the smallest valid value without loss of generality. Leveraging this classification of actions, we can bound the indicator for selecting a suboptimal action $A_{t}$:
Lemma 3 (Indicator decomposition lemma).
For every $0\leq\ell\leq\lfloor\log k\rfloor$, we have
$\phantom{{}={}}\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]$
$\leq\frac{2^{\ell+1}}{k_{A_{t}}}\sum_{i\notin A^{*}}\mathbb{1}\left[\left(\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{\left(\sqrt{2}\right)^{\ell}}{16}\cdot\frac{\Delta_{A_{t}}}{k_{A_{t}}}\right)\land i\in A_{t}\right].$
Note that the coefficient $\frac{2^{\ell+1}}{k_{A_{t}}}$ is critical to control the quantity of the summation. Then we need to solve challenge 3. We choose to give the actions in every $\mathcal{A}_{\ell}$ a regret upper bound. That is, for every $0\leq\ell\leq\lfloor\log k\rfloor$,
$\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]=O\left(\sqrt{kmT}\right).$
For any fixed $0\leq\ell\leq\lfloor\log k\rfloor$, we bound the regret caused by executing actions in $\mathcal{A}_{\ell}$. After applying Lemma 3, we recount the regret by switching the summation of base arms with the summation of actions that a certain arm is involved. To avoid redundant computations for regret, we could arrange the sequence of actions that maximizes the value.
Let each $\mathtt{arm}_{i}$ be contained in $N_{i,\ell}$ suboptimal actions $A_{i,\ell,1},\cdots,A_{i,\ell,N_{i,\ell}}$ such that $A_{i,\ell,n}\in\mathcal{A}_{\ell},\forall n\in[N_{i,\ell}]$. Denote $\Delta_{A_{i,\ell,n}}$ as $\Delta_{i,\ell,n}$ and $k_{A_{i,\ell,n}}$ as $k_{i,\ell,n}$. Without loss of generality, suppose that $\frac{\Delta_{i,\ell,N_{i,\ell}}}{k_{i,\ell,N_{i,\ell}}}\geq\frac{\Delta_{i,\ell,N_{i,\ell}-1}}{k_{i,\ell,N_{i,\ell}-1}}\geq\cdots\geq\frac{\Delta_{i,\ell,1}}{k_{i,\ell,1}}$. Then we have the following lemma:
Lemma 4.
For every $0\leq\ell\leq\lfloor\log k\rfloor$, we have
$\phantom{{}={}}\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\Delta_{A_{t}}\right]$
$\leq 1540\sum_{i\notin A^{*}}\Bigg(\sum_{n=2}^{N_{i,\ell}}\frac{k_{i,\ell,n}^{2}\ln\left(\frac{k^{2}}{\delta}\cdot\frac{\Delta_{i,\ell,n}^{2}}{k_{i,\ell,n}^{2}}\right)}{\Delta_{i,\ell,n}^{2}}\cdot$
$\phantom{{}={}}\left(\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}-\frac{\Delta_{i,\ell,n-1}}{k_{i,\ell,n-1}}\right)+\frac{k_{i,\ell,1}\ln\left(\frac{k^{2}}{\delta}\cdot\frac{\Delta_{i,\ell,1}^{2}}{k_{i,\ell,1}^{2}}\right)}{\Delta_{i,\ell,1}}\Bigg).$
By perceiving $\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}$ as the variable $x$, we could bound the first term by the integral of $f(x)=\frac{\ln\left(\frac{k^{2}x^{2}}{\delta}\right)}{x^{2}}$ over the relevant domain, while the second term by the maximum of $g(x)=\frac{\ln\left(\frac{k^{2}x^{2}}{\delta}\right)}{x}$ over its effective domain. Therefore we have
$\phantom{{}={}}\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{B}\right]\Delta_{A_{t}}\right]$
$=\sum_{\ell=0}^{\lfloor\log k\rfloor}\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\Delta_{A_{t}}\right]$
$=O\left((\log k)\sqrt{kmT}\right).$
Bring this back to Equation 1, we could obtain $\mathbf{E}\left[R(T)\right]=\mathbf{E}\left[\sum_{t=1}^{T}\Delta_{A_{t}}\right]=O\left((\log k)\sqrt{kmT}\right)$.
3.3 Extension to cascading feedback
In this section we introduce the cascading feedback (Kveton et al. 2015a, b; Li et al. 2016) and give our regret upper bound. In this model, when $A_{t}$ is chosen, playing this action means that the player reveals the outcomes of arms one by one following the sequence order until certain stopping condition is satisfied. The feedback is the outcomes of revealed arms and the reward is a function form of all arms in $A_{t}$. In particular, in the disjunctive form the player stops when the first 1 is revealed, or he reaches the end and gets all feedback. In the conjunctive form, the player stops when the first 0 is revealed or she reaches the end with all 1 outcomes.
Let $p_{A,t}$ be the the probability of observing all base arms of $A$ at round $t$. And denote $p^{*}=\min_{t\in T}\min_{A\in\mathcal{A}}p_{A,t}$, which is the minimal probability that an action could have all base arms observed over all time. Then we have the following regret bound:
Theorem 5.
For any stochastic combinatorial cascading feedback instance with $m$ arms, time horizon $T$, and the maximum number of arms that could be chosen in a round is $k$, then the expected regret of Algorithm 1 is $O\left(\frac{\log k}{p^{*}}\sqrt{kmT}\right)$ when $k\leq\frac{m}{2}$ and $O\left(\frac{m-k}{p^{*}}\sqrt{\log k\log(m-k)T}\right)$ when $k>\frac{m}{2}$.
The proof is given in Section A.2. In Section C.4, we also present comparative experiments evaluating the performance of CUCB and CMOSS under cascading feedback.
4 Experiments
Figure 1: Comparison of CMOSS (blue) with three baselines algorithms under semi-bandit feedback. Subplots (1)-(4), (9)-(12) use synthetic dataset; (5)-(8) use Yelp dataset. Subplots (1)-(8) show cumulative regret with $k=10,m=30$ and $k=20,m=30$; (9)(10) show ablation studies varying $k$ (fixed $m=30$), while (11)(12) varying $m$ (fixed $k=15$). Initial means of base arms fall within the range $[0,0.1]$, except for (2)(4)(6)(8), which use the range $[0.3,0.4]$.
In this section, we evaluate CMOSS on both synthetic and real-world datasets under semi-bandit feedback, benchmarking against three representative baselines: CUCB, EXP3.M, and HYBRID. In addition, we conduct ablation studies to investigate the influence of key parameters and assess the robustness of CMOSS under various configurations. All results are averaged over 10 independent runs of $100000$ rounds, including those in Appendix C. Further experimental details and additional ablation results are provided in Appendix C. Visual comparisons across all experiments are summarized in Figure 1.
For all experiments, we fix the confidence parameter of CMOSS to $\delta=0.00001$, which is of the same order of magnitude as the theoretical results $\frac{m(\log k)^{2}}{kT},k\leq\frac{m}{2}$ and $\frac{(m-k)^{2}\log k\log(m-k)}{k^{2}T},k>\frac{m}{2}$ (see Section 3.1 and Section A.1). For EXP3.M, we use a mixing coefficient $\gamma=0.01$, as it yields favorable empirical performance in our settings (see Section C.3). All rewards follow Bernoulli distributions: pulling arm $\mathtt{arm}_{i}$ yields reward $1$ with probability $\mu_{i}$, and 0 otherwise, where $\mu_{i}$ denotes the true mean (interpretable as the latent click-through probability in recommendation tasks).
4.1 Comparison with baselines
We evaluate CMOSS across both synthetic and real-world bandit environments to assess its effectiveness under diverse reward structures. (1)-(8) of Figure 1 report the cumulative regret over time with cardinality $k=10$ and $k=20$ and fixed number of base arms $m=30$.
In the synthetic setting, the initial click probability $\mu_{i}$ for each arm $\mathtt{arm}_{i}$ is independently drawn from a uniform distribution. We consider two regimes: a low-probability regime with $\mu_{i}\sim\mathtt{U}(0,0.1)$, and a high-probability regime with $\mu_{i}\sim\mathtt{U}(0.3,0.4)$. To further assess the practical effectiveness of our method, we evaluate it on the public Yelp dataset [^1] http://www.yelp.com/dataset˙challenge, a classical CMAB real-world dataset. Following (Li, Chen, and Leung 2019), we map user–item affinities to the target regimes $[0,0.1]$ and $[0.3,0.4]$ via linear rescaling. Full details are provided in Section C.1. The resulting values serve as the initial click probabilities of base arms in the two settings. For each regime, we independently sample $30$ arms.
The low-probability regime $[0,0.1]$ reflects typical conditions observed in real-world recommendation scenarios, where user interactions such as clicks occur relatively infrequently. This setting provides a realistic environment for evaluating algorithm performance. In contrast, the high-probability regime $[0.3,0.4]$ represents a more idealized scenario with higher click probabilities, allowing us to assess algorithm behavior in more optimistic conditions. By examining both regimes, we gain complementary insights into the robustness and effectiveness of the algorithms across a range of practical and theoretical contexts.
[…]
4.2 Ablation studies
To further evaluate the robustness and scalability of CMOSS, we conduct ablation studies by varying cardinality $k$ while fixing base arm size at $m=30$ and varying base arm size $m$ while fixing cardinality $k=15$ on synthetic dataset. As shown (3)(4) (varying $k$) and (7)(8) (varying $m$) of Figure 1, we report both cumulative regret and total runtime. Detailed metrics are summarized in Table 3. We also do the same ablation studies on Yelp dataset (see details in Section C.2).
Across both synthetic and Yelp datasets, CMOSS consistently achieves the lowest cumulative regret when varying either $k$ or $m$. Relative to CUCB, CMOSS reduces regret by 40–55% in most configurations, while requiring only a 10–30% increase in runtime. When compared with EXP3.M and HYBRID, CMOSS offers significant improvements: regret is typically 25–70% lower, and runtime is reduced by more than 65–85% across nearly all settings. The ablation studies on the Yelp dataset reveal the same performance patterns as those on the synthetic dataset, confirming that CMOSS remains robust and efficient across different domains and parameter regimes.
In addition to the main experiments, we include supplementary studies in Appendix C to strengthen the completeness and rigor of our evaluation. We first perform a parameter selection study for EXP3.M (see Section C.3) to validate the choice of its hyperparameter $\gamma=0.01$ in our setup. We then compare CMOSS and CUCB under a cascading feedback model (see Section C.4) to assess their performance under an alternative user interaction mechanism. In particular, CMOSS achieves a notable reduction in regret under cascading feedback, while incurring only a modest increase in runtime. These additional experiments support both the soundness of our experimental design and the robustness of CMOSS across varied settings.
These results indicate that CMOSS consistently outperforms benchmark algorithms in both regret and runtime efficiency across different datasets and varying values of $k$ and $m$, demonstrating robust effectiveness. This consistent performance highlights CMOSS’s suitability for practical recommendation and ranking tasks, demonstrating its effectiveness and reliability across varying selection sizes.
5 Conclusion and Future Work
In this paper, we study the stochastic CMAB problem and propose the CMOSS algorithm, which is not only computationally efficient but also has a regret of $O\big((\log k)\sqrt{kmT}\big)$ when $k\leq\frac{m}{2}$ and $O\big((m-k)\sqrt{\log k\log(m-k)T}\big)$ when $k>\frac{m}{2}$ under semi-bandit feedback. We then extend our analysis to cascading feedback. Comparative experiments with CUCB, EXP3.M and HYBRID under semi-bandit feedback show that CMOSS has at least one-way significant advantages in cumulative regret and computation cost over the other three algorithms, with the other one showing small differences.
Several promising directions for future research emerge from this work. First, tightening the $\frac{1}{p^{*}}$ factor under cascading feedback and other scenarios which contain probabilistically triggered arms, could yield a more general framework and facilitate the analysis of a broader range of problems including episodic reinforcement learning (Liu et al. 2024). Furthermore, recent refinements to the UCB algorithm (Lattimore 2015; Mukherjee et al. 2018), were shown to empirically outperform MOSS in terms of regret. Adapting these more sophisticated UCB variants to the CMAB setting presents a natural and compelling extension.
Appendix A Techinical Proofs of Section 3
A.1 Proofs of Section 3.2
For completeness and consistency of the paper, we first provide proofs of some basic lemmas (Lemma 7, Lemma 8, Lemma 9) that will be used later.
Lemma 6 (Hoeffding’s lemma, (Mitzenmacher and Upfal 2017)).
Let $X$ be a random variable with mean 0 and $a\leq X\leq b$ holds almost surely. Then for any $\lambda\in\mathbb{R}$,
$\mathbf{E}\left[e^{\lambda X}\right]\leq e^{\frac{\lambda^{2}(b-a)^{2}}{8}}.$
Lemma 7 (Generalization of Hoeffding’s inequality, (Mitzenmacher and Upfal 2017)).
Let $(Y_{t})_{t=1}^{n}$ be a sequence of independent variables in $[0,1]$ with means $(q_{t})_{t=1}^{n}$. let $X_{t}=Y_{t}-q_{t}$ and $S_{t}=\sum_{s=1}^{t}X_{s}$. Then for any $\varepsilon>0$,
$\mathbf{Pr}\left[\max_{t\in[n]}S_{t}\geq\varepsilon\right]\leq e^{-\frac{2\varepsilon^{2}}{n}}.$
Proof.
First we show that $(e^{\lambda S_{t}})^{n}_{t=1}$ is a submartingale for any $\lambda\in\mathbb{R}$. Let $\mathcal{F}_{t}=\sigma\left(X_{1},\dots,X_{t}\right)$. Note that $e^{\lambda x}$ is a convex function for all $\lambda\in\mathbb{R}$. Applying conditional Jensen’s inequality, for all $t\geq 1$:
$\mathbf{E}\left[e^{\lambda S_{t}}|\mathcal{F}_{t-1}\right]\geq e^{\lambda\mathbf{E}\left[S_{t}|\mathcal{F}_{t-1}\right]}=e^{\lambda S_{t-1}}.$
Therefore $(e^{\lambda S_{t}})^{n}_{t=1}$ is a submartingale with regard to $\left\{\mathcal{F}_{t}\right\}_{t\in[N]}$. From the Doob’s submartingale inequality, for any $\lambda>0$,
$\mathbf{Pr}\left[\underset{t\in[n]}{\max}S_{t}\geq\epsilon\right]=\mathbf{Pr}\left[\underset{t\in[n]}{\max}e^{\lambda S_{t}}\geq e^{\lambda\epsilon}\right]\leq\frac{\mathbf{E}\left[e^{\lambda S_{n}}\right]}{e^{\lambda\epsilon}}.$
According to Lemma 6, $\mathbf{E}\left[e^{\lambda X_{t}}\right]\leq e^{\frac{\lambda^{2}(b-a)^{2}}{8}}=e^{\frac{\lambda^{2}}{8}}$, then we have
$\mathbf{Pr}\left[\underset{t\in[n]}{\max}S_{t}\geq\epsilon\right]\leq\frac{\mathbf{E}\left[e^{\lambda S_{n}}\right]}{e^{\lambda\epsilon}}=\frac{\prod_{i=1}^{n}\mathbf{E}\left[e^{\lambda X_{i}}\right]}{e^{\lambda\epsilon}}\leq e^{\frac{n\lambda^{2}}{8}-\epsilon\lambda}.$
Set $\lambda=\frac{4\varepsilon}{n}$ to minimize $\frac{n\lambda^{2}}{8}-\epsilon\lambda$, we obtain
$\mathbf{Pr}\left[\underset{t\in[n]}{\max}S_{t}\geq\epsilon\right]\leq e^{-\frac{2\varepsilon^{2}}{n}}.$
Lemma 8 ((Lattimore and Szepesvári 2020)).
Define $\zeta_{i}=\max\left(0,\mu_{i}-\min\limits_{t\leq T}\left(\hat{\mu}_{i,t}+\rho_{i,t}\right)\right)$. For any $\varepsilon>0$,
$\mathbf{Pr}\left[\zeta_{i}\geq\varepsilon\right]\leq\frac{4\delta}{\varepsilon^{2}},\quad\mathbf{E}\left[\zeta_{i}\right]\leq 4\sqrt{\delta}.$
Proof.
$\mathbf{Pr}\left[\zeta_{i}\geq\varepsilon\right]=0$ when $\epsilon\geq\mu_{i}$. Therefore we only need to prove the case when $0<\epsilon<\mu_{i}$. First we have:
$\mathbf{Pr}\left[\zeta_{i}\geq\epsilon\right]\leq\mathbf{Pr}\left[\max_{s\leq T}\left(\mu_{i}-\left(\hat{\mu}_{i,s}+\sqrt{\frac{1}{s}\ln^{+}(\frac{1}{\delta s}})\right)\right)\geq\varepsilon\right].$
Without loss of generality, let $T=2^{L}$, $S_{t}=\sum_{s=1}^{t}\left(\mu_{i}-X_{i}^{(s)}\right)$. It’s easy to verify that $(e^{\lambda S_{t}})^{T}_{t=1}$ is a submartingale for any $\lambda\in\mathbb{R}$. We have:
$\phantom{{}={}}\mathbf{Pr}\left[\max_{s\leq T}\left(\mu_{i}-\left(\hat{\mu}_{i,s}+\sqrt{\frac{1}{s}\ln^{+}(\frac{1}{\delta s}})\right)\right)\geq\varepsilon\right]$
$\leq\sum_{j=1}^{L}\mathbf{Pr}\left[\max_{2^{j-1}\leq s\leq 2^{j}}s\left(\mu_{i}-\hat{\mu}_{i,s}\right)\geq 2^{j-1}\varepsilon+\sqrt{2^{j-1}\ln^{+}(\frac{1}{\delta\cdot 2^{j}}})\right]$
$\leq\sum_{j=1}^{L}\mathbf{Pr}\left[\max_{1\leq s\leq 2^{j}}S_{s}\geq 2^{j-1}\varepsilon+\sqrt{2^{j-1}\ln^{+}(\frac{1}{\delta\cdot 2^{j}}})\right]$
$\leq\sum_{j=1}^{L}\exp\left(-\frac{\left(2^{j-1}\varepsilon+\sqrt{2^{j-1}\ln^{+}(\frac{1}{\delta\cdot 2^{j}}})\right)^{2}}{2^{j-1}}\right)$ (2)
$\leq\delta\sum_{j=1}^{L}2^{j}e^{-2^{j-1}\varepsilon^{2}}$
$\leq\frac{\delta}{e\varepsilon^{2}}+\delta\int_{0}^{+\infty}2^{x+1}e^{-2^{x}\varepsilon^{2}}dx$ (3)
$=\frac{\delta}{e\varepsilon^{2}}+\frac{2\delta}{\ln 2\cdot\varepsilon^{2}e^{\varepsilon}}\leq\frac{4\delta}{\varepsilon^{2}},$
where Equation 2 follows from Lemma 7. And Equation 3 follows from
$\sum_{x=a}^{b}f(x)\leq\max_{x\in[a,b]}f(x)+\int_{a}^{b}f(x)dx$
when $f(x)$ is unimodal. Therefore, we have $\mathbf{Pr}\left[\zeta_{i}\geq\varepsilon\right]\leq\frac{4\delta}{\varepsilon^{2}}$. And
$\mathbf{E}\left[\zeta_{i}\right]=\int_{0}^{1}\mathbf{Pr}\left[\zeta_{i}\geq t\right]dt\leq\int_{0}^{1}\min\left\{1,\frac{4\delta}{t^{2}}\right\}dt\leq 4\sqrt{\delta}.$
Lemma 9 ((Lattimore and Szepesvári 2020)).
Let $X_{1},\dots,X_{n}$ be a sequence of independent random variables in $[0,1]$. Let $\hat{\mu}_{t}=\frac{1}{t}\sum_{s=1}^{t}X_{s}$. For a fixed $a>1$, $\varepsilon\in(0,1)$, define
$\kappa=\sum_{t=1}^{T}\mathbb{1}\left[\hat{\mu}_{t}+\sqrt{\frac{a}{t}}-\mu\geq\varepsilon\right],$
then for any integer $T>1$, we have $\mathbf{E}\left[\kappa\right]\leq\frac{3a}{\varepsilon^{2}}$.
Proof.
Denote $S_{t}=\sum_{s=1}^{t}(X_{s}-\mu)$, so $S_{t}$ is a martingale. We have:
$\mathbf{E}\left[\kappa\right]$ $=\sum_{t=1}^{n}\mathbf{Pr}\left[S_{t}\geq t\varepsilon-\sqrt{at}\right]$
$=\sum_{t=1}^{\lfloor a/\varepsilon^{2}\rfloor}\left(1-\mathbf{Pr}\left[S_{t}<t\varepsilon-\sqrt{at}\right]\right)+\sum_{t=\lfloor a/\varepsilon^{2}\rfloor+1}^{n}\mathbf{Pr}\left[S_{t}\geq t\varepsilon-\sqrt{at}\right]$
$\leq\frac{a}{\varepsilon^{2}}+\sum_{t=\lfloor a/\varepsilon^{2}\rfloor+1}^{n}\exp\left(-\frac{2\left(t\varepsilon-\sqrt{at}\right)^{2}}{t}\right)$
$=\frac{a}{\varepsilon^{2}}+\sum_{t=\lfloor a/\varepsilon^{2}\rfloor+1}^{n}e^{-2\left(\sqrt{t}\varepsilon-\sqrt{a}\right)^{2}}$
$\leq\frac{a}{\varepsilon^{2}}+\int_{\frac{a}{\varepsilon^{2}}}^{+\infty}e^{-2\left(\sqrt{t}\varepsilon-\sqrt{a}\right)^{2}}dt$
$=\frac{a}{\varepsilon^{2}}+\frac{\sqrt{2\pi}}{2}\cdot\frac{\sqrt{a}}{\varepsilon^{2}}+\frac{1}{2\varepsilon^{2}}\leq\frac{3a}{\varepsilon^{2}}.$
We begin to give a thorough proof of Theorem 1. When Algorithm 1 chooses $A_{t}$, for the arms in $A_{t}\setminus A^{*}$ and arms in $A^{*}\setminus A_{t}$, we have
$\min\limits_{i\in A_{t}\setminus A^{*}}\left(\hat{\mu}_{i,t}+\rho_{i,t}\right)\geq\max\limits_{j\in A^{*}\setminus A_{t}}\left(\hat{\mu}_{j,t}+\rho_{j,t}\right)\geq\frac{1}{\left|A^{*}\setminus A_{t}\right|}\sum_{j\in A^{*}\setminus A_{t}}\left(\hat{\mu}_{j,t}+\rho_{j,t}\right).$
Then for every $i\in A_{t}\setminus A^{*}$, we have
$\hat{\mu}_{i,t}+\rho_{i,t}\geq\frac{1}{\left|A^{*}\setminus A_{t}\right|}\min\limits_{s\leq T}\left(\sum_{j\in A^{*}\setminus A_{t}}\left(\hat{\mu}_{j,s}+\rho_{j,s}\right)\right)\geq\frac{1}{\left|A^{*}\setminus A_{t}\right|}\sum_{j\in A^{*}\setminus A_{t}}\min\limits_{s\leq T}\left(\hat{\mu}_{j,s}+\rho_{j,s}\right).$
According to the definition, $\zeta_{j}\geq\mu_{j}-\min\limits_{s\leq T}\left(\hat{\mu}_{j,s}+\rho_{i,s}\right),\forall j\in A^{*}$. Therefore
$\hat{\mu}_{i,t}+\rho_{i,t}$ $\geq\frac{1}{\left|A^{*}\setminus A_{t}\right|}\sum_{j\in A^{*}\setminus A_{t}}\left(\mu_{j}-\zeta_{j}\right)\geq\frac{1}{\left|A^{*}\setminus A_{t}\right|}\left(\sum_{j\in A^{*}\setminus A_{t}}\mu_{j}-\sum_{j\in A^{*}}\zeta_{j}\right)$
Since $\Delta_{A_{t}}=\sum_{j\in A^{*}\setminus A_{t}}\mu_{j}-\sum_{i\in A_{t}\setminus A^{*}}\mu_{i}$ and $\Delta_{A_{t}}>8\sum_{j\in A^{*}}\zeta_{j}$, then
$\hat{\mu}_{i,t}+\rho_{i,t}\geq\frac{1}{\left|A^{*}\setminus A_{t}\right|}\left(\sum_{j\in A^{*}\setminus A_{t}}\mu_{j}-\sum_{j\in A^{*}}\zeta_{j}\right)\geq\frac{1}{\left|A^{*}\setminus A_{t}\right|}\left(\sum_{i\in A_{t}\setminus A^{*}}\mu_{i}+\frac{7\Delta_{A_{t}}}{8}\right).$
Remember that we denote $\left(\mu^{A}_{i}\right)_{i\in[k_{A}]}$ as the mean rewards of the arms in $A\setminus A^{*}$ where $\mu^{A}_{k_{A}}\geq\mu^{A}_{k_{A}-1}\geq\cdots\geq\mu^{A}_{1}$ and $\overline{\mu^{A}}=\frac{\sum_{i\in[k_{A}]}\mu^{A}_{i}}{k_{A}}$. Define $D_{A}=\left\{i\in[k_{A}]\;\middle|\;\mu^{A}_{i}\leq\overline{\mu^{A}}+\frac{3\Delta_{A}}{4k_{A}}\right\}$ as the set of arms in $A\setminus A^{*}$ whose mean reward is no more than $\overline{\mu^{A}}+\frac{3\Delta_{A}}{4k_{A}}$. Then we have Lemma 2:
Lemma 10 (Lemma 2 restated).
For every suboptimal action $A\in\mathcal{A}$, if $\left|D_{A}\right|\leq\lfloor\frac{k_{A}}{2}\rfloor$, then there exists $1\leq\ell_{A}\leq\lfloor\log k_{A}\rfloor$ such that
$\overline{\mu^{A}}-\mu^{A}_{\lfloor\frac{k_{A}}{2^{\ell_{A}}}\rfloor}\geq\frac{\left(\sqrt{2}\right)^{\ell_{A}}}{16}\cdot\frac{\Delta_{A}}{k_{A}}.$
Proof.
Since $\left|D_{A}\right|\leq\lfloor\frac{k_{A}}{2}\rfloor$, we have $\mu^{A}_{\lfloor\frac{k_{A}}{2}\rfloor+1}>\overline{\mu^{A}}+\frac{3\Delta_{A}}{4k_{A}}$. From the definition of $\overline{\mu^{A}}$, we obtain
$\sum_{i=1}^{\lfloor\frac{k_{A}}{2}\rfloor}\left(\overline{\mu^{A}}-\mu^{A}_{i}\right)$ $=\sum_{i=\lfloor\frac{k_{A}}{2}\rfloor+1}^{k_{A}}\left(\mu^{A}_{i}-\overline{\mu^{A}}\right)$
$\geq\sum_{i=\lfloor\frac{k_{A}}{2}\rfloor+1}^{k_{A}}\left(\mu^{A}_{\lfloor\frac{k_{A}}{2}\rfloor+1}-\overline{\mu^{A}}\right)$
$>\left(k_{A}-\lfloor\frac{k_{A}}{2}\rfloor\right)\cdot\frac{3\Delta_{A}}{4k_{A}}\geq\frac{3\Delta_{A}}{8}.$ (4)
We prove the lemma by contradiction, Assume that for all $1\leq\ell\leq\lfloor\log k_{A}\rfloor$, $\overline{\mu^{A}}-\mu^{A}_{\lfloor\frac{k_{A}}{2^{\ell}}\rfloor}<\frac{\left(\sqrt{2}\right)^{\ell}}{16}\cdot\frac{\Delta_{A}}{k_{A}}$. So for all $\lfloor\frac{k_{A}}{2^{\ell+1}}\rfloor+1\leq i\leq\lfloor\frac{k_{A}}{2^{\ell}}\rfloor$, we have
$\overline{\mu^{A}}-\mu^{A}_{i}\leq\overline{\mu^{A}}-\mu^{A}_{\lfloor\frac{k_{A}}{2^{\ell+1}}\rfloor}<\frac{\left(\sqrt{2}\right)^{\ell+1}}{16}\cdot\frac{\Delta_{A}}{k_{A}}.$
Then we bound $\sum_{i=1}^{\lfloor\frac{k_{A}}{2}\rfloor}\left(\overline{\mu^{A}}-\mu^{A}_{i}\right)$ as follows:
$\sum_{i=1}^{\lfloor\frac{k_{A}}{2}\rfloor}\left(\overline{\mu^{A}}-\mu^{A}_{i}\right)$ $=\sum_{j=1}^{\lfloor\log k_{A}\rfloor}\sum_{i=\lfloor\frac{k_{A}}{2^{j+1}}\rfloor+1}^{\lfloor\frac{k_{A}}{2^{j}}\rfloor}\left(\overline{\mu^{A}}-\mu^{A}_{i}\right)$
$<\sum_{j=1}^{\lfloor\log k_{A}\rfloor}\left(\lfloor\frac{k_{A}}{2^{j}}\rfloor-\lfloor\frac{k_{A}}{2^{j+1}}\rfloor\right)\cdot\frac{\left(\sqrt{2}\right)^{j+1}}{16}\cdot\frac{\Delta_{A}}{k_{A}}$
$\leq\sum_{j=1}^{\lfloor\log k_{A}\rfloor}\left(\frac{k_{A}}{2^{j+1}}+\frac{1}{2}\right)\cdot\frac{\left(\sqrt{2}\right)^{j+1}}{16}\cdot\frac{\Delta_{A}}{k_{A}}$
$=\sum_{j=1}^{\lfloor\log k_{A}\rfloor}\frac{\Delta_{A}}{16\left(\sqrt{2}\right)^{j+1}}+\sum_{j=1}^{\lfloor\log k_{A}\rfloor}\frac{\left(\sqrt{2}\right)^{j+1}\Delta_{A}}{32k_{A}}$
$<\frac{\Delta_{A}}{8}+\frac{\left(\sqrt{2}+1\right)\Delta_{A}}{16\sqrt{k_{A}}}<\frac{3\Delta_{A}}{8},$
which contradicts Equation 4. ∎
Now we divide the proof into two cases where $k\leq\frac{m}{2}$ and $k>\frac{m}{2}$.
The case where $k\leq\frac{m}{2}$
We focus on proving the third term $\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{B}\right]\Delta_{A_{t}}\right]$, where
$\mathcal{A}_{B}=\left\{A\in\mathcal{A}\;\middle|\;\Delta_{A}>\max\left\{8\sum_{j\in A^{*}}\zeta_{j},(\log k)\sqrt{\frac{ekm}{T}}\right\}\right\}.$
From Lemma 2, before the algorithm starts, we could divide the action set $\mathcal{A}$ into $1+\lfloor\log k\rfloor$ exhaustive and mutually exclusive sets $\left(\mathcal{A}_{\ell}\right)_{0\leq\ell\leq\lfloor\log k\rfloor}$ since $1\leq k_{A}\leq k$, where
$\mathcal{A}_{0}=\left\{A\in\mathcal{A}\;\middle|\;\left|D_{A}\right|>\lfloor\frac{k_{A}}{2}\rfloor\right\},\mathcal{A}_{\ell}=\left\{A\in\mathcal{A}\;\middle|\;\left|D_{A}\right|\leq\lfloor\frac{k_{A}}{2}\rfloor\land\ell_{A}=\ell\right\}.$
Leveraging this classification of actions, we can bound the indicator for selecting a suboptimal action $A_{t}$ through this lemma:
Lemma 11 (Lemma 3 restated).
For every $0\leq\ell\leq\lfloor\log k\rfloor$, we have
$\phantom{{}={}}\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\leq\frac{2^{\ell+1}}{k_{A_{t}}}\sum_{i\notin A^{*}}\mathbb{1}\left[\left(\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{\left(\sqrt{2}\right)^{\ell}}{16}\cdot\frac{\Delta_{A_{t}}}{k_{A_{t}}}\right)\land i\in A_{t}\right].$
Proof.
By the definition of $D_{A}$, we know that for every $A_{t}\in\mathcal{A}_{0}$, there are at least $\lfloor\frac{k_{A_{t}}}{2}\rfloor+1$ arms ($\mathtt{arm}_{i}$ e.g.) in $A_{t}\setminus A^{*}$ satisfies
$\hat{\mu}_{i,t}+\rho_{i,t}\geq\overline{\mu^{A_{t}}}+\frac{7\Delta_{A_{t}}}{8k_{A_{t}}}\geq\mu_{i}+\frac{\Delta_{A_{t}}}{8k_{A_{t}}}\geq\mu_{i}+\frac{\Delta_{A_{t}}}{16k_{A_{t}}}.$
Therefore we have
$\mathbb{1}\left[A_{t}\in\mathcal{A}_{0}\cap\mathcal{A}_{B}\right]\leq\frac{2}{k_{A_{t}}}\sum_{i\notin A^{*}}\mathbb{1}\left[\left(\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{\Delta_{A_{t}}}{16k_{A_{t}}}\right)\land i\in A_{t}\right].$
For every $A_{t}\in\mathcal{A}_{\ell},1\leq\ell\leq\lfloor\log k\rfloor$, there are at least $\lfloor\frac{k_{A_{t}}}{2^{\ell}}\rfloor$ arms ($\mathtt{arm}_{i}$ e.g.) in $A_{t}\setminus A^{*}$ satisfies
$\hat{\mu}_{i,t}+\rho_{i,t}\geq\overline{\mu^{A_{t}}}+\frac{7\Delta_{A_{t}}}{8k_{A_{t}}}\geq\mu_{i}+\frac{\left(\sqrt{2}\right)^{\ell}}{16}\cdot\frac{\Delta_{A_{t}}}{k_{A_{t}}}.$
Similarly we have
$\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]$ $\leq\frac{1}{\lfloor\frac{k_{A_{t}}}{2^{\ell}}\rfloor}\sum_{i\notin A^{*}}\mathbb{1}\left[\left(\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{\left(\sqrt{2}\right)^{\ell}}{16}\cdot\frac{\Delta_{A_{t}}}{k_{A_{t}}}\right)\land i\in A_{t}\right]$
$\leq\frac{2^{\ell+1}}{k_{A_{t}}}\sum_{i\notin A^{*}}\mathbb{1}\left[\left(\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{\left(\sqrt{2}\right)^{\ell}}{16}\cdot\frac{\Delta_{A_{t}}}{k_{A_{t}}}\right)\land i\in A_{t}\right].$
In summary, for every $0\leq\ell\leq\lfloor\log k\rfloor$,
$\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\leq\frac{2^{\ell+1}}{k_{A_{t}}}\sum_{i\notin A^{*}}\mathbb{1}\left[\left(\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{\left(\sqrt{2}\right)^{\ell}}{16}\cdot\frac{\Delta_{A_{t}}}{k_{A_{t}}}\right)\land i\in A_{t}\right].$
Lemma 3 quantitatively decouples the constrains for $A_{t}$ into every arm in $A_{t}\setminus A^{*}$. Next we give the actions in every $\mathcal{A}_{\ell}$ a regret upper bound. That is, for every $0\leq\ell\leq\lfloor\log k\rfloor$, we want to prove
$\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]=O\left(\sqrt{kmT}\right).$
Recall that each $\mathtt{arm}_{i}$ is contained in $N_{i,\ell}$ suboptimal actions $A_{i,\ell,1},\cdots,A_{i,\ell,N_{i,\ell}}$ such that $A_{i,\ell,n}\in\mathcal{A}_{\ell},\forall n\in[N_{i,\ell}]$. Denote $\Delta_{A_{i,\ell,n}}$ as $\Delta_{i,\ell,n}$ and $k_{A_{i,\ell,n}}$ as $k_{i,\ell,n}$. Without loss of generality, suppose that $\frac{\Delta_{i,\ell,N_{i,\ell}}}{k_{i,\ell,N_{i,\ell}}}\geq\frac{\Delta_{i,\ell,N_{i,\ell}-1}}{k_{i,\ell,N_{i,\ell}-1}}\geq\cdots\geq\frac{\Delta_{i,\ell,1}}{k_{i,\ell,1}}$. Then we have:
$\phantom{{}={}}\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]$
$\leq\sum_{t=1}^{T}\mathbf{E}\left[\sum_{i\notin A^{*}}\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{A_{t}}}{k_{A_{t}}}\right]\cdot\frac{2^{\ell+1}\Delta_{A_{t}}}{k_{A_{t}}}\;\middle|\;i\in A_{t}\right]$
$=\sum_{i\notin A^{*}}\sum_{t=1}^{T}\sum_{n=1}^{N_{i,\ell}}\mathbf{E}\left[\mathbb{1}\left[A_{t}=A_{i,\ell,n}\right]\cdot\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\cdot\frac{2^{\ell+1}\Delta_{i,\ell,n}}{k_{i,\ell,n}}\;\middle|\;i\in A_{t}\right]$
$=2^{\ell+1}\sum_{i\notin A^{*}}\mathbf{E}\left[\sum_{t=1}^{T}\sum_{n=1}^{N_{i,\ell}}\mathbb{1}\left[A_{t}=A_{i,\ell,n}\right]\cdot\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\;\middle|\;i\in A_{t}\right],$ (5)
where the first inequality is from Lemma 3. To avoid redundant computations for regret at Equation 5, we could arrange the sequence of actions that maximizes the value. Moreover, applying Lemma 9, we have the following lemma:
Lemma 12 (Lemma 4 restated).
For every $0\leq\ell\leq\lfloor\log k\rfloor$, we have
$\phantom{{}={}}\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\Delta_{A_{t}}\right]$
$\leq 1540\sum_{i\notin A^{*}}\left(\sum_{n=2}^{N_{i,\ell}}\frac{k_{i,\ell,n}^{2}\ln\left(\frac{k^{2}}{\delta}\cdot\frac{\Delta_{i,\ell,n}^{2}}{k_{i,\ell,n}^{2}}\right)}{\Delta_{i,\ell,n}^{2}}\left(\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}-\frac{\Delta_{i,\ell,n-1}}{k_{i,\ell,n-1}}\right)+\frac{k_{i,\ell,1}\ln\left(\frac{k^{2}}{\delta}\cdot\frac{\Delta_{i,\ell,1}^{2}}{k_{i,\ell,1}^{2}}\right)}{\Delta_{i,\ell,1}}\right).$
Proof.
Since $\frac{\Delta_{i,\ell,N_{i,\ell}}}{k_{i,\ell,N_{i,\ell}}}\geq\frac{\Delta_{i,\ell,N_{i,\ell}-1}}{k_{i,\ell,N_{i,\ell}-1}}\geq\cdots\geq\frac{\Delta_{i,\ell,1}}{k_{i,\ell,1}}$, when $\hat{\mu}_{i,t}+\rho_{i,t}-\mu_{i}\geq\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,N_{i,\ell}}}{k_{i,\ell,N_{i,\ell}}}$, we should choose $A_{t}=A_{i,\ell,N_{i,\ell}}$. And when $\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,N_{i,\ell-1}}}{k_{i,\ell,N_{i,\ell-1}}}\leq\hat{\mu}_{i,t}+\rho_{i,t}-\mu_{i}<\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,N_{i,\ell}}}{k_{i,\ell,N_{i,\ell}}}$, we should choose $A_{t}=A_{i,\ell,N_{i,\ell}-1}$, and so on. By choosing $A_{1},\cdots A_{T}$ to maximize the value of
$\sum_{t=1}^{T}\sum_{n=1}^{N_{i,\ell}}\mathbb{1}\left[A_{t}=A_{i,\ell,n}\right]\cdot\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}},$
we obtain that
$\phantom{{}={}}\sum_{t=1}^{T}\sum_{n=1}^{N_{i,\ell}}\mathbb{1}\left[A_{t}=A_{i,\ell,n}\right]\cdot\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}$
$\leq\sum_{t=1}^{T}\bigg(\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,N_{i,\ell}}}{k_{i,\ell,N_{i,\ell}}}\right]\cdot\frac{\Delta_{i,\ell,N_{i,\ell}}}{k_{i,\ell,N_{i,\ell}}}$
$\phantom{{}={}}+\sum_{n=1}^{N_{i,\ell}-1}\mathbb{1}\left[\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\leq\hat{\mu}_{i,t}+\rho_{i,t}-\mu_{i}<\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n+1}}{k_{i,\ell,n+1}}\right]\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\bigg)$
$=\sum_{t=1}^{T}\bigg(\sum_{n=2}^{N_{i,\ell}}\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\left(\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}-\frac{\Delta_{i,\ell,n-1}}{k_{i,\ell,n-1}}\right)$
$\phantom{{}={}}+\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,1}}{k_{i,\ell,1}}\right]\cdot\frac{\Delta_{i,\ell,1}}{k_{i,\ell,1}}\bigg),$ (6)
where Equation 6 is due to $\mathbb{1}\left[x\leq y<z\right]=\mathbb{1}\left[y\geq x\right]-\mathbb{1}\left[y\geq z\right]$ for any $y$ if $x<z$.
From Lemma 9, for every $\Delta_{i,\ell,n},n\in[N_{i,\ell}]$, we bound
$\mathbf{E}\left[\sum_{t=1}^{T}\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\;\middle|\;i\in A_{t}\right]$
as follows:
$\phantom{{}={}}\mathbf{E}\left[\sum_{t=1}^{T}\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\;\middle|\;i\in A_{t}\right]$
$\leq\mathbf{E}\left[\sum_{T_{i,t}=1}^{T}\mathbb{1}\left[\frac{\sum_{s=1}^{T_{i,t}}X_{i}^{(s)}}{T_{i,t}}+\sqrt{\frac{\ln^{+}\left(\frac{1}{\delta T_{i,t}}\right)}{T_{i,t}}}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\right]$
$\leq\frac{1}{\Delta_{i,\ell,n}^{2}}+\mathbf{E}\left[\sum_{T_{i,t}>\frac{1}{\Delta_{i,\ell,n}^{2}}}^{T}\mathbb{1}\left[\frac{\sum_{s=1}^{T_{i,t}}X_{i}^{(s)}}{{T_{i,t}}}+\sqrt{\frac{\ln\left(\frac{\Delta_{i,\ell,n}^{2}}{\delta}\right)}{T_{i,t}}}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\right]$
$\leq\frac{1}{\Delta_{i,\ell,n}^{2}}+\frac{768k_{i,\ell,n}^{2}\ln\left(\frac{\Delta_{i,\ell,n}^{2}}{\delta}\right)}{2^{\ell}\Delta_{i,\ell,n}^{2}}\leq\frac{770k_{i,\ell,n}^{2}\ln\left(\frac{\Delta_{i,\ell,n}^{2}}{\delta}\right)}{2^{\ell}\Delta_{i,\ell,n}^{2}}.$
Combining this with Equation 5 and Equation 6, we have
$\phantom{{}={}}\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]$
$\leq 2^{\ell+1}\sum_{i\notin A^{*}}\mathbf{E}\Bigg[\sum_{t=1}^{T}\sum_{n=1}^{N_{i,\ell}}\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\Bigg|\,i\in A_{t}\Bigg]$
$\leq 2^{\ell+1}\sum_{i\notin A^{*}}\mathbf{E}\Bigg[\sum_{t=1}^{T}\bigg(\sum_{n=2}^{N_{i,\ell}}\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\left(\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}-\frac{\Delta_{i,\ell,n-1}}{k_{i,\ell,n-1}}\right)$
$\phantom{{}={}}\phantom{{}={}}+\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,1}}{k_{i,\ell,1}}\right]\cdot\frac{\Delta_{i,\ell,1}}{k_{i,\ell,1}}\bigg)\Bigg|\,i\in A_{t}\Bigg]$
$=2^{\ell+1}\sum_{i\notin A^{*}}\bigg(\sum_{n=2}^{N_{i,\ell}}\mathbf{E}\Bigg[\sum_{t=1}^{T}\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\Bigg|\,i\in A_{t}\Bigg]\left(\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}-\frac{\Delta_{i,\ell,n-1}}{k_{i,\ell,n-1}}\right)$
$\phantom{{}={}}\phantom{{}={}}+\mathbf{E}\Bigg[\sum_{t=1}^{T}\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,1}}{k_{i,\ell,1}}\right]\Bigg|\,i\in A_{t}\Bigg]\cdot\frac{\Delta_{i,\ell,1}}{k_{i,\ell,1}}\bigg)$
$\leq 770\cdot 2^{\ell+1}\sum_{i\notin A^{*}}\left(\sum_{n=2}^{N_{i,\ell}}\frac{k_{i,\ell,n}^{2}\ln\left(\frac{\Delta_{i,\ell,n}^{2}}{\delta}\right)}{2^{\ell}\Delta_{i,\ell,n}^{2}}\left(\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}-\frac{\Delta_{i,\ell,n-1}}{k_{i,\ell,n-1}}\right)+\frac{k_{i,\ell,1}^{2}\ln\left(\frac{\Delta_{i,\ell,1}^{2}}{\delta}\right)}{2^{\ell}\Delta_{i,\ell,1}^{2}}\cdot\frac{\Delta_{i,\ell,1}}{k_{i,\ell,1}}\right)$
$\leq 1540\sum_{i\notin A^{*}}\left(\sum_{n=2}^{N_{i,\ell}}\frac{k_{i,\ell,n}^{2}\ln\left(\frac{k^{2}}{\delta}\cdot\frac{\Delta_{i,\ell,n}^{2}}{k_{i,\ell,n}^{2}}\right)}{\Delta_{i,\ell,n}^{2}}\left(\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}-\frac{\Delta_{i,\ell,n-1}}{k_{i,\ell,n-1}}\right)+\frac{k_{i,\ell,1}\ln\left(\frac{k^{2}}{\delta}\cdot\frac{\Delta_{i,\ell,1}^{2}}{k_{i,\ell,1}^{2}}\right)}{\Delta_{i,\ell,1}}\right).$
Perceiving $\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}$ as the variable $x$, we could bound the first term by the integral of $f(x)=\frac{\ln\left(\frac{k^{2}x^{2}}{\delta}\right)}{x^{2}}$ over $x>\frac{\Delta_{i,\ell,1}}{k_{i,\ell,1}}>(\log k)\sqrt{\frac{em}{kT}}$, while the second term is bounded by the maximum of $g(x)=\frac{\ln\left(\frac{k^{2}x^{2}}{\delta}\right)}{x}$ over $x>(\log k)\sqrt{\frac{em}{kT}}$. $f(x)$ has been plotted in Figure 2. It is easy to verify that $f(x)$ is monotonically decreasing and positive when $x>\sqrt{\frac{e\delta}{k^{2}}}=(\log k)\sqrt{\frac{em}{k^{3}T}}$. Additionally we have $\int f(x)dx=\frac{\ln\frac{\delta}{e^{2}k^{2}}-2\ln x}{x}$. Thus the summation could be bounded by the integral of $f(x)$ on $((\log k)\sqrt{\frac{em}{kT}},\infty)$. Also $g(x)$ is monotonically decreasing when $x>(\log k)\sqrt{\frac{em}{kT}}$, which gives an upper bound for the second term $\frac{k_{i,\ell,1}\ln\left(\frac{k^{2}}{\delta}\cdot\frac{\Delta_{i,\ell,1}^{2}}{k_{i,\ell,1}^{2}}\right)}{\Delta_{i,\ell,1}}$. Thus we have
$\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\right]\Delta_{A_{t}}\right]$ $\leq 1540\sum_{i\notin A^{*}}\left(\int_{(\log k)\sqrt{\frac{em}{kT}}}^{\infty}f(x)dx+\frac{\ln ek^{2}}{(\log k)\sqrt{\frac{em}{kT}}}\right)$
$=1540\sum_{i\notin A^{*}}\left(\frac{\ln\frac{\delta}{e^{2}k^{2}}-2\ln x}{x}\Big|_{(\log k)\sqrt{\frac{em}{kT}}}^{\infty}+\frac{\ln ek^{2}}{\log k}\sqrt{\frac{kT}{em}}\right)$
$=1540\sum_{i\notin A^{*}}\left(\frac{\ln e^{4}k^{4}}{\log k}\sqrt{\frac{kT}{em}}\right)=O\left(\sqrt{kmT}\right).$
[…]
Figure 2: The plot of $f(x)=\frac{\ln\left(\frac{k^{2}x^{2}}{\delta}\right)}{x^{2}}$.
Therefore we have
$\sum_{t=1}^{T}\mathbf{E}\left[\left[A_{t}\in\mathcal{A}_{B}\right]\Delta_{A_{t}}\right]$ $=\sum_{\ell=0}^{\lfloor\log k\rfloor}\sum_{t=1}^{T}\mathbf{E}\left[\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\Delta_{A_{t}}\right]$
$=O\left((\log k)\sqrt{kmT}\right).$
Bring this back to Equation 1, we could obtain $\mathbf{E}\left[R(T)\right]=\mathbf{E}\left[\sum_{t=1}^{T}\Delta_{A_{t}}\right]=O\left((\log k)\sqrt{kmT}\right)$.
The case where $k>\frac{m}{2}$
In this case, we need to prove that the regret of CMOSS is $O\left((m-k)\sqrt{\log k\log(m-k)T}\right)$. Here $\delta$ is optimized to be $\frac{(m-k)^{2}\log k\log(m-k)}{k^{2}T}$. Now we let $\mathcal{A}_{B}=\left\{A\in\mathcal{A}\;\middle|\;\Delta_{A}>\max\left\{8\sum_{j=1}^{k}\zeta_{j},(m-k)\sqrt{\frac{e\log k\log(m-k)}{T}}\right\}\right\}$. And we decompose the overall regret as follows:
$\sum_{t=1}^{T}\mathbf{E}\left[\Delta_{A_{t}}\right]$ $=\sum_{t=1}^{T}\mathbf{E}\left[\left[\Delta_{A_{t}}\leq(m-k)\sqrt{\frac{e\log k\log(m-k)}{T}}\right]\Delta_{A_{t}}\right]+\sum_{t=1}^{T}\mathbf{E}\left[\left[\Delta_{A_{t}}\leq 8\sum_{j=1}^{k}\zeta_{j}\right]\Delta_{A_{t}}\right]$
$+\sum_{t=1}^{T}\mathbf{E}\left[\left[\Delta_{A_{t}}>\max\left\{8\sum_{j=1}^{k}\zeta_{j},(m-k)\sqrt{\frac{e\log k\log(m-k)}{T}}\right\}\right]\Delta_{A_{t}}\right]$
$\leq(m-k)\sqrt{e\log k\log(m-k)T}+8\sum_{j=1}^{k}\mathbf{E}\left[\zeta_{j}\right]\cdot T+\sum_{t=1}^{T}\mathbf{E}\left[\left[A_{t}\in\mathcal{A}_{B}\right]\Delta_{A_{t}}\right].$ (7)
From Lemma 8, the second term could be bounded: $8\sum_{j=1}^{k}\mathbf{E}\left[\zeta_{j}\right]\cdot T\leq 8k\sqrt{\delta}\cdot T=8(m-k)\sqrt{\log k\log(m-k)T}$.
Since now the range of $k_{A}$ changes to $1\leq k_{A}\leq m-k$, from Lemma 2, we could divide the action set $\mathcal{A}$ into $1+\lfloor\log(m-k)\rfloor$ exhaustive and mutually exclusive sets $\left(\mathcal{A}_{\ell}\right)_{0\leq\ell\leq\lfloor\log(m-k)\rfloor}$, where
$\mathcal{A}_{0}=\left\{A\in\mathcal{A}\;\middle|\;\left|D_{A}\right|>\lfloor\frac{k_{A}}{2}\rfloor\right\},\mathcal{A}_{\ell}=\left\{A\in\mathcal{A}\;\middle|\;\left|D_{A}\right|\leq\lfloor\frac{k_{A}}{2}\rfloor\land\ell_{A}=\ell\right\}.$
Furthermore, we could similarly prove Lemma 3 in this case: for every $0\leq\ell\leq\lfloor\log(m-k)\rfloor$,
$\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\leq\frac{2^{\ell+1}}{k_{A_{t}}}\sum_{i\notin A^{*}}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{\left(\sqrt{2}\right)^{\ell}}{16}\cdot\frac{\Delta_{A_{t}}}{k_{A_{t}}}\land i\in A_{t}\right].$ (8)
Next we give the actions in every $\mathcal{A}_{\ell}$ a regret upper bound. That is, for every $0\leq\ell\leq\lfloor\log(m-k)\rfloor$, we need to prove
$\sum_{t=1}^{T}\mathbf{E}\left[\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]=O\left((m-k)\sqrt{\frac{T\log k}{\log(m-k)}}\right).$
Following the same process of the proof of Lemma 4, by slightly changing the upper bound of $k_{A}$ from $k$ to mkm-k, we have that, for every $0\leq\ell\leq\lfloor\log(m-k)\rfloor$,
$\phantom{{}={}}\sum_{t=1}^{T}\mathbf{E}\left[\left[A_{t}\in\mathcal{A}_{\ell}\right]\Delta_{A_{t}}\right]$
$\leq 1540\sum_{i\notin A^{*}}\left(\sum_{n=2}^{N_{i,\ell}}\frac{k_{i,\ell,n}^{2}\ln\left(\frac{(m-k)^{2}}{\delta}\cdot\frac{\Delta_{i,\ell,n}^{2}}{k_{i,\ell,n}^{2}}\right)}{\Delta_{i,\ell,n}^{2}}\left(\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}-\frac{\Delta_{i,\ell,n-1}}{k_{i,\ell,n-1}}\right)+\frac{k_{i,\ell,1}\ln\left(\frac{(m-k)^{2}}{\delta}\cdot\frac{\Delta_{i,\ell,1}^{2}}{k_{i,\ell,1}^{2}}\right)}{\Delta_{i,\ell,1}}\right).$
Next we consider the properties of the function $f(x)=\frac{\ln\left(\frac{(m-k)^{2}x^{2}}{\delta}\right)}{x^{2}},x>\frac{\Delta_{i,\ell,1}}{k_{i,\ell,1}}>\sqrt{\frac{e\log k\log(m-k)}{T}}=\frac{k\sqrt{e\delta}}{m-k}$. It is easy to verify that $f(x)$ is monotonically decreasing and positive when $x>\sqrt{\frac{e\delta}{(m-k)^{2}}}=\frac{\sqrt{e\delta}}{m-k}$. Additionally we have $\int f(x)dx=\frac{\ln\frac{\delta}{e^{2}(m-k)^{2}}-2\ln x}{x}$. Also $g(x)=\frac{\ln\left(\frac{(m-k)^{2}x^{2}}{\delta}\right)}{x}$ is monotonically decreasing when $x>\frac{k\sqrt{e\delta}}{m-k}$. Thus the summation could be bounded by the integral of $f(x)$ on $(\frac{k\sqrt{e\delta}}{m-k},\infty)$
$\sum_{t=1}^{T}\mathbf{E}\left[\left[A_{t}\in\mathcal{A}_{\ell}\right]\Delta_{A_{t}}\right]$ $\leq 1540\sum_{i\notin S^{*}}\left(\int_{\frac{k\sqrt{e\delta}}{m-k}}^{\infty}f(x)dx+\frac{\ln ek^{2}}{\frac{k\sqrt{e\delta}}{m-k}}\right)$
$\leq 1540\sum_{i\notin S^{*}}\left(\frac{\ln\frac{\delta}{e^{2}(m-k)^{2}}-2\ln x}{x}\Big|_{\frac{k\sqrt{e\delta}}{m-k}}^{\infty}+\ln ek^{2}\sqrt{\frac{T}{e\log k\log(m-k)}}\right)$
$=1540\sum_{i\notin S^{*}}\left(\ln e^{4}k^{4}\sqrt{\frac{T}{e\log k\log(m-k)}}\right)$
$=O\left((m-k)\sqrt{\frac{T\log k}{\log(m-k)}}\right).$
Applying this lemma, we have
$\sum_{t=1}^{T}\mathbf{E}\left[\left[A_{t}\in\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]$ $=\sum_{\ell=0}^{\lfloor\log(m-k)\rfloor}\sum_{t=1}^{T}\mathbf{E}\left[\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]$
$=O\left((m-k)\sqrt{\log k\log(m-k)T}\right).$
Bring this back to Equation 7, we could obtain that the regret of Algorithm 1 when $k>\frac{m}{2}$ is $O\left((m-k)\sqrt{\log k\log(m-k)T}\right)$.
A.2 Proofs of Section 3.3
We take the disjunctive cascading bandits for example. In this case $r(A_{t},\bm{\mu})=\prod_{i\in A_{t}}(1-\mu_{i})$. Since $\mathcal{A}$ contains all subsets of $[m]$, $A^{*}$ consists of the the $k$ arms with the highest expected means. Denote $A^{*}=\left\{j_{1},\cdots,j_{k}\right\}$ where $\mu_{j_{1}}\leq\mu_{j_{2}}\leq\cdots\leq\mu_{j_{k}}$ and $A_{t}=\left\{i_{1},\cdots,i_{k}\right\}$ where $\mu_{i_{1}}\leq\cdots\leq\mu_{i_{k}}$. Thus
$\left|r(A^{*},\bm{\mu})-r(A_{t},\bm{\mu}))\right|$ $=\left|\prod_{n=1}^{k}(1-\mu_{j_{n}})-\prod_{n=1}^{k}(1-\mu_{i_{n}})\right|$
$\leq\sum_{l=0}^{k-1}\left|\prod_{n=1}^{l}(1-\mu_{i_{n}})\prod_{n=l+1}^{k}(1-\mu_{j_{n}})-\prod_{n=1}^{l+1}(1-\mu_{i_{n}})\prod_{n=l+2}^{k}(1-\mu_{j_{n}})\right|$
$=\sum_{l=0}^{k-1}\left(\prod_{n=1}^{l}(1-\mu_{i_{n}})\cdot\prod_{n=l+2}^{k}(1-\mu_{j_{n}})\cdot\left|\mu_{j_{l+1}}-\mu_{i_{l+1}}\right|\right)$
$\leq\sum_{n=1}^{k}\left|\mu_{j_{n}}-\mu_{i_{n}}\right|=\sum_{n=1}^{k}\mu_{j_{n}}-\sum_{n=1}^{k}\mu_{i_{n}}.$
Let $p_{A,t}$ be the the probability of observing all base arms of $A$ at round $t$. And denote $p^{*}=\min_{t\in T}\min_{A\in\mathcal{A}}p_{A,t}$. Therefore in each round $\left[\tau_{t}=A_{t}\right]\geq p^{*}$. Let $\mathcal{F}_{t}$ be the history before the learning agent chooses action at time $t$. Thus $\mathcal{F}_{t}$ contains feedback information at all time $s<t$. We could bound the regret as follows
$\mathbf{E}\left[R(T)\right]$ $=\mathbf{E}\left[\mathbf{E}\left[\sum_{t=1}^{T}\Delta_{A_{t}}\;\middle|\;\mathcal{F}_{t}\right]\right]$
$=\mathbf{E}\left[\mathbf{E}\left[\sum_{t=1}^{T}\Delta_{A_{t}}\cdot\mathbf{E}\left[\frac{1}{p_{A_{t},t}}\left[\tau_{t}=A_{t}\right]\;\middle|\;A_{t}\right]\;\middle|\;\mathcal{F}_{t}\right]\right]$
$\leq\frac{1}{p^{*}}\mathbf{E}\left[\mathbf{E}\left[\sum_{t=1}^{T}\Delta_{A_{t}}\left[\tau_{t}=A_{t}\right]\;\middle|\;\mathcal{F}_{t}\right]\right]$
$=\begin{cases}O\left(\frac{\log k}{p^{*}}\sqrt{kmT}\right)&,k\leq\frac{m}{2}\\ O\left(\frac{m-k}{p^{*}}\sqrt{\log k\log(m-k)T}\right)&,k>\frac{m}{2}\end{cases}$ (9)
where Equation 9 applies Theorem 1 to this case.
For the conjunctive case, $r(A_{t},\bm{\mu})=\prod_{i\in A_{t}}\mu_{i}$. The rest analysis follows the same pattern as the disjunctive case.
Appendix B Details of CUCB, EXP3.M and HYBRID
In this section we present the complete formulations of the CUCB algorithm (Chen, Wang, and Yuan 2013; Kveton et al. 2015c; Liu et al. 2024; Wang and Chen 2017), the EXP3.M algorithm (Uchiya, Nakamura, and Kudo 2010) and the HYBRID algorithm (Zimmert, Luo, and Wei 2019).
B.1 The CUCB algorithm
Algorithm 2 Combinatorial Upper Confidence Bound (CUCB)
Input:$m,k$.
1: For each arm $i$, $T_{i}\leftarrow 0$ $\triangleright$ maintain the total number of times arm $i$ is played so far
2: For each arm $i$, $\hat{\mu}_{i}\leftarrow 1$ $\triangleright$ maintain the empirical mean of $\mathtt{arm}_{i}$
3: for $t=1,2,3,\ldots$ do
4:   For each arm $i\in[m]$, $\rho_{i}\leftarrow\sqrt{\frac{3\ln t}{2T_{i}}}$ $\triangleright$ the confidence radius, $\rho_{i}=+\infty$ if $T_{i}=0$
5:   For each arm $i\in[m]$, $\bar{\mu}_{i}\leftarrow\min\left\{\hat{\mu}_{i}+\rho_{i},1\right\},\bar{}\bm{\mu}\leftarrow(\bar{\mu}_{1},\cdots,\bar{\mu}_{m})$
6:   $A_{t}\leftarrow\mathop{\arg\max}_{A\in\mathcal{A}}r(A,\bar{}\bm{\mu})$
7:   Pull $A_{t}$ of base arms and obtain the triggered set $\tau_{t}$. Observe feedback $X_{i}^{(t)}$ for all $i\in\tau_{t}$
8:   For every $i\in\tau_{t}$, update $T_{i}$ and $\hat{\mu}_{i}$: $T_{i}\leftarrow T_{i}+1$, $\hat{\mu}_{i}\leftarrow\hat{\mu}_{i}+(X_{i}^{(t)}-\hat{\mu}_{i})/T_{i}$
The regret bound of the CUCB algorithm is $O\left(\sqrt{kmT\log T}\right)$ under semi-bandit feedback. A detailed analysis of the regret is provided in (Kveton et al. 2015c).
B.2 The EXP3.M algorithm
Algorithm 3 EXP3.M
Input:mixing coefficient $\gamma\in(0,1]$
1: $\omega_{i}(1)=1$ for $i=1,\cdots,m$ $\triangleright$ Initialization
2: for $t=1,2,3,\ldots$ do
3:   if $\max_{j\in[m]}\omega_{j}(t)\geq\left(\frac{1}{k}-\frac{\gamma}{m}\right)\sum_{i=1}^{m}\frac{\omega_{i}(t)}{1-\gamma}$ then
4:    Decide $\alpha_{t}$ so as to satisfy $\frac{\alpha_{t}}{\sum_{\omega_{i}(t)\geq\alpha_{t}}\alpha_{t}+\sum_{\omega_{i}(t)<\alpha_{t}}\omega_{i}(t)}=\left(\frac{1}{k}-\frac{\gamma}{m}\right)\cdot\frac{1}{1-\gamma}$
5:    Set $S_{0}(t)=\left\{i\;\middle|\;\omega_{i}(t)\geq\alpha_{t}\right\}$ and $\omega_{i}^{\prime}(t)=\alpha_{t}$ for $i\in S_{0}(t)$
6:   else
7:    Set $S_{0}(t)=\emptyset$
8:   Set $\omega_{i}^{\prime}(t)=\omega_{i}(t)$ for $i\in[m]\setminus S_{0}(t)$
9:   Set $p_{i}(t)=k\left(\left(1-\gamma\right)\frac{\omega_{i}^{\prime}(t)}{\sum_{j=1}^{m}\omega_{j}^{\prime}(t)}+\frac{\gamma}{m}\right)$ for $i\in[m]$
10:   Set $S(t)=\mathbf{DepRound}\left(k,(p_{1}(t),p_{2}(t),\cdots,p_{m}(t))\right)$
11:   Receive rewards $x_{i}(t)$ for $i\in S(t)$
12:   for $i=1,\cdots,m$ do
13:    Set $\hat{x}_{i}(t)=\begin{cases}\frac{x_{i}(t)}{p_{i}(t)},&i\in S(t)\\ 0,&o.w.\end{cases}$
14:    Set $\omega_{i}(t+1)=\begin{cases}\omega_{i}(t)\exp\left(\frac{k\gamma\hat{x}_{i}(t)}{m}\right),&i\notin S_{0}(t)\\ \omega_{i}(t),&o.w.\end{cases}$
In Algorithm 3, if all $\omega_{i}(t)$ are less than $\left(\frac{1}{k}-\frac{\gamma}{m}\right)\sum_{i=1}^{m}\frac{\omega_{i}(t)}{1-\gamma}$, then every $p_{i}(t)$ calculated at Line 9 of 9 is less than $1$ and $S_{0}(t)=\emptyset$. Otherwise it chooses a appropriate threshold $\alpha_{t}$ that the temporal weight $\omega_{i}^{\prime}(t)$ is set to $\alpha_{t}$ for $i\in S_{0}(t)$ and $\omega_{i}(t)$ for $i\notin S_{0}(t)$. Thus $p_{i}(t)=1$ for $i\in S_{0}(t)$, which means that $S_{0}(t)\subseteq S(t)$. And $\omega_{i}(t)$ is not updated for $i\in S_{0}(t)$ since it is large to some extent.
The algorithm $\mathbf{DepRound}$ is designed as follows:
Algorithm 4 $\mathbf{DepRound}$
Input: Natural number $k$, $\left(p_{1},\cdots,p_{m}\right)$ with $\sum_{i=1}^{m}p_{i}=k,0\leq p_{i}\leq 1,\forall i\in[m]$
1: while there exists $i\in[m]$ with $0<p_{i}<1$ do
2:   Choose distinct $i$ and $j$ with $0<p_{i}<1$ and $0<p_{j}<1$
3:   Set $\alpha=\min\left\{1-p_{i},p_{j}\right\}$, $\beta=\min\left\{p_{i},1-p_{j}\right\}$
4:   Update $p_{i},p_{j}$ as $(p_{i},p_{j})=\begin{cases}(p_{i}+\alpha,p_{j}-\alpha),&w.p.\;\frac{\beta}{\alpha+\beta}\\ (p_{i}-\beta,p_{j}+\beta),&w.p.\;\frac{\alpha}{\alpha+\beta}\end{cases}$
output $\left\{i\;\middle|\;p_{i}=1,i\in[m]\right\}$
In Algorithm 4, $(p_{1},\cdots,p_{m})$ is probabilistically updated until all the components are 0 or $1$ while keeping the condition that $\sum_{i=1}^{m}p_{i}=k$. Therefore if there exists $i\in[m]$ with $0<p_{i}<1$, then there must be some $j\neq i$ with $0<p_{j}<1$ otherwise the summation $\sum_{i=1}^{m}p_{i}$ is not an integer. Inside the while-loop, the iteration lasts at most $m$ times because at least one of $p_{i}$ and $p_{j}$ becomes 0 or $1$ each time. The regret bound of the EXP3.M algorithm is $O\left(\sqrt{kmT\log\left(\frac{m}{k}\right)}\right)$ and the detailed analysis is provided in (Uchiya, Nakamura, and Kudo 2010).
B.3 The HYBRID algorithm
Algorithm 5 FTRL with hybrid regularizer for semi-bandits
Input: $0<\gamma\leq 1$, sampling scheme $P$
1: $\hat{L}_{0}=(0,\ldots,0),\eta_{t}=1/\sqrt{t}$
2: for $t=1,2,\dots$ do
3:   Compute $x_{t}=\mathop{\arg\min}\limits_{x\in\operatorname*{\operatorname{Conv}}(\mathcal{A})}\big\langle x,\hat{L}_{t-1}\big\rangle+\eta_{t}^{-1}\Psi(x)$
4:   Sample $A_{t}\sim P(x_{t})$
5:   Observe $o_{t}=A_{t}\circ\ell_{t}$
6:   Construct estimator $\hat{\ell}_{t},\;\forall i:\;\hat{\ell}_{ti}=\frac{(o_{ti}+1)\mathbb{I}_{t}(i)}{x_{ti}}-1$
7:   Update $\hat{L}_{t}=\hat{L}_{t-1}+\hat{\ell}_{t}$
Here $\operatorname*{\operatorname{Conv}}(\mathcal{A})$ is the convex hull of $\mathcal{A}$. And the hybrid regularizer
$\Psi(x)=\sum_{i=1}^{m}-\sqrt{x_{i}}+\gamma(1-x_{i})\log(1-x_{i})$ (10)
with a parameter $0<\gamma\leq 1$ to be chosen as follows
$\gamma=\begin{cases}1&\mbox{ if }k\leq\frac{m}{2}\\ \min\{1,1/\sqrt{\log(m/(m-k))}\}&\mbox{ otherwise.}\end{cases}$
The regret bound of the HYBRID algorithm is $O\left(\sqrt{kmT}\right)$ when $k\leq\frac{m}{2}$ and $O\left((m-k)\sqrt{\log\left(\frac{m}{m-k}\right)T}\right)$ when $k>\frac{m}{2}$. And the detailed analysis is provided in (Zimmert, Luo, and Wei 2019).
Appendix C Experiments
The experiments are performed on a Dell Inspiration 16 PLUS 7620 laptop equipped with a 12th Generation Intel(R) Core(TM) i7-12700H CPU running at 2.30 GHz, alongside 16 GB of RAM. The operating system utilized is Windows 11, and the Python version employed is 3.11.7.
C.1 Comparison with baselines and beyond
Synthetic dataset
As described in the main text, in the synthetic setting, we simulate a combinatorial bandit problem with $m=30$ base arms and cardinality $k=10$ and $k=20$. The initial value for each arm is independently drawn from uniform distributions $\mathtt{U}(0,0.1)$ and $\mathtt{U}(0.3,0.4)$. Each algorithm is run for $100000$ rounds, and all results are averaged over $10$ independent runs.
[…]
Real-world dataset
To simulate a real-world bandit environment, we use the public Yelp dataset [^2] http://www.yelp.com/dataset˙challenge, a classical dataset in the CMAB literature. Following the setup in (Li, Chen, and Leung 2019), we construct a filtered subset consisting of 1500 users and 10000 items from the original dataset. Each user and item is represented as a $50$-dimensional real-valued feature vector, normalized to unit $\ell_{2}$ norm. For each user–item pair $(u,a)$, we compute an affinity score as the dot product $\mathbf{u}^{\top}\mathbf{a}$, where $\mathbf{u},\mathbf{a}\in\mathbb{R}^{50}$.
To match the synthetic regimes used in our experiments, the raw affinity scores are min-max normalized and linearly rescaled into two ranges: a low-probability regime $[0,0.1]$ and a high-probability regime $[0.3,0.4]$. Specifically, let $\min$ and $\max$ denote the minimum and maximum dot products across all user–item pairs. The scores are transformed as follows:
$\text{score}_{\text{low}}(u,a)$ $=0.1\cdot\frac{\mathbf{u}^{\top}\mathbf{a}-\min}{\max-\min},$ (11)
$\text{score}_{\text{high}}(u,a)$ $=0.3+0.1\cdot\frac{\mathbf{u}^{\top}\mathbf{a}-\min}{\max-\min}.$ (12)
The resulting values are interpreted as the initial click probabilities for base arms. For each regime, we independently sample 30 such values as base arms using a fixed random seed (2026) to ensure reproducibility.
[…]
Table 2: Semi-bandit feedback: cumulative regret and runtime (in seconds) of algorithms under two initial base arm value ranges: $[0,0.1]$ and $[0.3,0.4]$, evaluated on both synthetic and Yelp datasets and fixed $m=30$.
$\star$ indicates our proposed method (CMOSS).
Metric Algorithm Range $[0,0.1]$ Range $[0.3,0.4]$
Synthetic Yelp Synthetic Yelp
$k=10$ $k=20$ $k=10$ $k=20$ $k=10$ $k=20$ $k=10$ $k=20$
[…]
CUCB 4000.132 2348.979 4407.244 1916.620 3945.550 2361.077 4050.910 2214.263
EXP3.M 3523.649 1535.670 3094.234 1154.182 4009.885 2800.222 3555.937 2621.060
HYBRID 5858.858 3233.744 5197.661 2561.223 4706.725 2538.204 4067.236 2411.709
[…]
CUCB 15.155 27.816 11.422 21.108 10.724 20.391 11.739 23.882
EXP3.M 104.291 92.774 79.239 78.895 79.592 87.363 95.136 110.756
HYBRID 121.126 111.946 108.540 105.515 95.413 76.987 143.529 163.490
C.2 Ablation studies and beyond
We provide the complete results of our ablation study evaluating the impact of two key parameters: the cardinality $k$ of the super arm and the total number of base arms $m$. We consider two controlled experiments with initial base arms in $[0,0.1]$:
  • Varying $k$: $k\in\{5,10,15,20,25\}$, with $m=30$ fixed.
  • Varying $m$: $m\in\{20,25,30,35,40\}$, with $k=15$ fixed.
On the synthetic dataset ($\mathtt{U}(0,0.1)$), CMOSS consistently achieves the lowest cumulative regret across all configurations. When varying $k$, CMOSS outperforms CUCB, EXP3.M, and HYBRID by up to 52.7%, 55.7%, and 67.4%, respectively. When varying $m$, the gains become even more pronounced—CMOSS reduces regret by as much as 53.3% over CUCB, 40.6% over EXP3.M, and 68.7% over HYBRID. In terms of runtime, CMOSS remains close to CUCB while being substantially faster than EXP3.M and HYBRID across all settings. Moreover, CMOSS exhibits stable runtime as $m$ increases, demonstrating strong scalability. On the Yelp dataset (rescaled to $[0,0.1]$), we observe consistent trends. When varying $k$, CMOSS achieves regret reductions of up to 51.0% over CUCB, 39.4% over EXP3.M, and 70.0% over HYBRID. When varying $m$, CMOSS again demonstrates clear advantages—outperforming CUCB by up to 46.9%, 23.1%, and 61.5%, respectively. Runtime comparisons further confirm CMOSS’s efficiency: it remains close to CUCB while being substantially faster than both EXP3.M and HYBRID. Additionally, CMOSS maintains stable runtime as either $k$ or $m$ increases, showcasing its practical scalability on real-world data. Complete numerical results are reported in Table 3 and Table 4.
Table 3: Ablation study of semi-bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=15$ fixed) on synthetic dataset. Results correspond to subplots (3)(4)(7)(8) in Figure 1.
$\star$ indicates our proposed method (CMOSS).
Metric Algorithm Varying $k$ (Synthetic) Varying $m$ (Synthetic)
$k=5$ $k=10$ $k=15$ $k=20$ $k=25$ $m=20$ $m=25$ $m=30$ $m=35$ $m=40$
[…]
CUCB 4429.727 3446.924 3184.634 2578.753 1131.007 1296.234 2182.487 2791.642 4208.573 5586.506
EXP3.M 4736.340 3271.894 2162.514 1462.691 659.203 836.489 1500.403 2327.747 3305.964 4369.808
HYBRID 5596.137 5123.518 4676.580 3456.369 1517.378 1918.722 3376.063 4285.878 6283.876 8184.837
[…]
CUCB 6.574 11.654 16.397 21.287 26.991 16.218 21.619 25.239 18.210 69.932
EXP3.M 76.208 120.789 107.957 116.241 106.603 61.540 96.724 124.403 129.031 447.195
HYBRID 125.178 135.696 108.276 91.548 88.490 63.963 96.495 147.526 172.044 245.882
Table 4: Ablation study of semi-bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=15$ fixed) on Yelp dataset.
$\star$ indicates our proposed method (CMOSS).
Metric Algorithm Varying $k$ (Yelp) Varying $m$ (Yelp)
$k=5$ $k=10$ $k=15$ $k=20$ $k=25$ $m=20$ $m=25$ $m=30$ $m=35$ $m=40$
[…]
CUCB 4020.879 4208.577 3074.333 2175.718 1191.837 1162.516 2450.990 3299.031 3559.703 4558.815
EXP3.M 4021.209 2988.275 2025.406 1216.121 663.273 658.334 1409.273 2088.690 2424.766 3387.021
HYBRID 4969.380 5078.731 3960.924 2890.771 1944.711 1548.021 3338.523 4347.092 4913.297 5471.507
[…]
CUCB 7.448 13.630 19.281 25.876 31.900 16.236 24.373 23.632 28.706 71.921
EXP3.M 91.734 106.868 127.099 166.739 86.960 69.510 95.093 132.767 109.091 435.470
HYBRID 119.877 125.436 104.920 118.024 155.868 78.343 105.469 156.025 167.520 379.948
C.3 Parameter selection for EXP3.M
The performance of EXP3.M in semi-bandits is highly sensitive to the choice of its mixing coefficient $\gamma$, which controls the balance between exploration and exploitation. To determine an appropriate setting for our environment, we conduct an ablation over $\gamma\in\{0.001,0.01,0.1\}$ with cardinality $k\in\{10,20\}$. Each configuration is run for a horizon of $T=100000$ and averaged over 10 independent repetitions.
As shown in Table 5 and Figure 3, $\gamma=0.01$ consistently achieves the lowest cumulative regret across both initialization regimes. Under $\mathtt{U}(0,0.1)$, it reduces regret by $83.9\%$ and $80.3\%$ relative to $\gamma=0.001$ and $\gamma=0.1$ for $k=10$, and by $84.1\%$ and $44.8\%$ for $k=20$. A similar pattern holds under $\mathtt{U}(0.3,0.4)$: $\gamma=0.01$ improves over $\gamma=0.001$ and $\gamma=0.1$ by $75.6\%$ and $81.3\%$ for $k=10$, and by $67.3\%$ and $89.3\%$ for $k=20$. Based on these findings, we adopt $\gamma=0.01$ as the default parameter for all experiments involving EXP3.M, as it strikes the most favorable balance between exploration and exploitation in our stochastic environment.
Figure 3: Ablation study of EXP3.M with varying mixing coefficient $\gamma\in\{0.001,0.01,0.1\}$.
Table 5: Ablation study of EXP3.M under different exploration parameters $\gamma$ on the synthetic dataset. Results are shown under two base-arm initial value distributions: (1) $\mathtt{U}(0,0.1)$; (2) $\mathtt{U}(0.3,0.4)$.
Metric Algorithm Range $[0,0.1]$ Range $[0.3,0.4]$
$k=10$ $k=20$ $k=10$ $k=20$
$\gamma=0.001$ $\gamma=0.01$ $\gamma=0.1$ $\gamma=0.001$ $\gamma=0.01$ $\gamma=0.1$ $\gamma=0.001$ $\gamma=0.01$ $\gamma=0.1$ $\gamma=0.001$ $\gamma=0.01$ $\gamma=0.1$
Regret EXP3.M 21475.680 3457.137 17551.325 8570.484 1357.281 2459.607 17459.760 4255.882 22816.466 9514.021 3108.351 29068.987
Runtime (sec) EXP3.M 113.997 111.016 139.714 119.251 92.074 99.780 95.312 89.501 90.747 109.813 86.766 62.930
C.4 Cascading bandit
Furthermore, we simulate a disjunctive cascading bandit problem with $m=30$ base arms and cardinality $k=10$ and $k=20$. In each round, the algorithm recommends a list of $k$ items, and the user examines them sequentially—either in descending or ascending order of attraction probability—clicking on the first attractive item (if any) and ignoring the remaining ones. If none of the items is deemed attractive, the round results in no feedback and is treated as a no-op. The experimental configurations for both datasets (including synthetic and Yelp) and ablation studies in this section follow the same setup as detailed in Section 4 and Appendix C. Particularly, all experiments presented hereafter adopt this disjunctive feedback model. Visual comparisons across all experiments are summarized in Figure 4 (descending) and Figure 5 (ascending).
In addition, we do not include the configuration with $k=20,m=30$ under the $[0.3,0.4]$ regime. This is because, under the cascading setting, the regret gap becomes extremely small in this regime: with such a large slate size relative to the attraction range, the user almost always clicks on at least one item with high probability, making the optimal expected reward very close to 1 (as implied by the Bernoulli click model). As a result, different algorithms exhibit nearly indistinguishable performance, and the comparison becomes uninformative.
Descending bandit
[…]
Figure 4: Comparison of CMOSS (blue) with CUCB algorithms under cascading (descending) feedback. Subplots (1)(2)(5)(6) show cumulative regret with fixed $k=10$, $m=30$; (3)(4) show ablation studies varying $k$ (fixed $m=30$), while (7)(8) varying $m$ (fixed $k=15$). Subplots (1)(2)(3)(4)(7)(8) use synthetic dataset; (5)(6) use Yelp dataset. Initial means of base arms fall within the range $[0,0.1]$, except for (2)(6), which use the range $[0.3,0.4]$.
Table 6: Descending bandit feedback: cumulative regret and runtime (in seconds) of algorithms under two initial base arm value ranges: $[0,0.1]$ and $[0.3,0.4]$, evaluated on both synthetic and Yelp datasets and fixed $k=10,m=30$. Results correspond to subplots (1)(2)(5)(6) in Figure 4.
$\star$ indicates our proposed method (CMOSS).
Metric Algorithm Range $[0,0.1]$ Range $[0.3,0.4]$
Synthetic Yelp Synthetic Yelp
[…]
CUCB 2762.497 3214.506 250.376 161.587
[…]
CUCB 12.777 16.246 3.328 8.337
[…]
Table 7: Ablation study of descending bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=15$ fixed) on synthetic dataset within the range $[0,0.1]$. Results correspond to subplots (3)(4)(7)(8) in Figure 4.
$\star$ indicates our proposed method (CMOSS).
Metric Algorithm Varying $k$ (Synthetic) Varying $m$ (Synthetic)
$k=5$ $k=10$ $k=15$ $k=20$ $k=25$ $m=20$ $m=25$ $m=30$ $m=35$ $m=40$
[…]
CUCB 4429.727 3446.924 3184.634 2578.753 1131.007 916.646 1391.695 1952.693 2273.459 2751.108
[…]
CUCB 6.574 11.654 16.397 21.287 26.991 24.877 16.767 16.434 14.984 14.915
Table 8: Ablation study of descending bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=15$ fixed) on Yelp dataset within the range $[0,0.1]$.
$\star$ indicates our proposed method (CMOSS).
Metric Algorithm Varying $k$ (Yelp) Varying $m$ (Yelp)
$k=5$ $k=10$ $k=15$ $k=20$ $k=25$ $m=20$ $m=25$ $m=30$ $m=35$ $m=40$
[…]
CUCB 3236.333 2602.221 2150.017 1607.918 919.645 674.154 1290.947 2385.077 2659.719 3752.076
[…]
CUCB 8.215 13.148 16.720 19.717 22.924 25.212 20.810 17.956 17.513 15.967
Ascending bandit
[…]
Figure 5: Comparison of CMOSS (blue) with CUCB algorithms under cascading (ascending) feedback. Subplots (1)(2)(5)(6) show cumulative regret with fixed $k=10$, $m=30$; (3)(4) show ablation studies varying $k$ (fixed $m=30$), while (7)(8) varying $m$ (fixed $k=15$). Subplots (1)(2)(3)(4)(7)(8) use synthetic dataset; (5)(6) use Yelp dataset. Initial means of base arms fall within the range $[0,0.1]$, except for (2)(6), which use the range $[0.3,0.4]$.
Table 9: Ascending bandit feedback: cumulative regret and runtime (in seconds) of algorithms under two initial base arm value ranges: $[0,0.1]$ and $[0.3,0.4]$, evaluated on both synthetic and Yelp datasets, with fixed $k=10,m=30$. Results correspond to subplots (1)(2)(5)(6) in Figure 5.
$\star$ indicates our proposed method (CMOSS).
Metric Algorithm Range $[0,0.1]$ Range $[0.3,0.4]$
Synthetic Yelp Synthetic Yelp
[…]
CUCB 2083.085 3010.909 3945.550 4050.910
[…]
CUCB 13.395 18.039 10.724 11.739
[…]
Table 10: Ablation study of ascending bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=15$ fixed) on synthetic dataset. Results correspond to subplots (3)(4)(7)(8) in Figure 5.
$\star$ indicates our proposed method (CMOSS).
Metric Algorithm Varying $k$ (Synthetic) Varying $m$ (Synthetic)
$k=5$ $k=10$ $k=15$ $k=20$ $k=25$ $m=20$ $m=25$ $m=30$ $m=35$ $m=40$
[…]
CUCB 4429.727 3446.924 3184.634 2578.753 1131.007 436.651 755.159 1041.633 1283.360 1752.840
[…]
CUCB 6.574 11.654 16.397 21.287 26.991 25.692 19.102 16.068 16.542 16.562
Table 11: Ablation study of ascending bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=15$ fixed) on Yelp dataset.
$\star$ indicates our proposed method (CMOSS).
Metric Algorithm Varying $k$ (Yelp) Varying $m$ (Yelp)
$k=5$ $k=10$ $k=15$ $k=20$ $k=25$ $m=20$ $m=25$ $m=30$ $m=35$ $m=40$
[…]
CUCB 3958.232 2224.800 1618.416 915.447 402.428 492.410 1059.454 1615.740 2240.731 2589.184
[…]
CUCB 11.738 13.694 21.308 29.020 31.540 26.714 21.569 19.727 18.980 14.424
Overall, these results demonstrate that CMOSS consistently achieves superior performance compared to CUCB across both descending- and ascending-order cascading bandit models. It significantly reduces cumulative regret across all datasets and configurations of cardinality $k$ and number of base arms $m$, while maintaining competitive runtime efficiency. Notably, CMOSS scales well with increasing $m$, often running faster or with only moderate overhead compared to CUCB, especially in larger-scale settings. This robust performance across diverse scenarios highlights CMOSS’s practicality and reliability, making it a strong candidate for real-world recommendation and ranking applications where both accuracy and computational cost are critical.